### v1

# Regularization in Data-driven Predictive Control: A Convex Relaxation Perspective

Abstract
This paper explores the role of regularization in data-driven predictive control (DDPC) through the lens of convex relaxation, in line with [1]. Using a bi-level optimization framework, we model system identification as an inner problem and predictive control as an outer problem. Within this framework, we show that several regularized DDPC formulations, including $\ell_{1}$-norm penalties, projection-based regularizers, and a newly introduced causality-based regularizer, can be viewed as convex relaxations of their respective bi-level problems. This perspective clarifies the conceptual links between direct and indirect data-driven control and highlights how regularization implicitly enforces system identification. We further propose an optimality-based variant, O-DDPC, which approximately solves the inner problem with all identification constraints via an iterative algorithm. Numerical experiments demonstrate that O-DDPC outperforms existing regularized DDPC by reducing both bias and variance errors. These results indicate that further benefits may be obtained by applying system identification techniques to pre-process the trajectory library in nonlinear settings. Overall, our analysis contributes to a unified convex relaxation view of regularization in DDPC and sheds light on its strong empirical performance beyond linear time-invariant systems.
keywords:
Data-driven Control, Bi-level optimization, Convex approximation
thanks: This work is supported by NSF ECCS-2154650, CMMI-2320697, CAREER 2340713, and an Early Career Faculty Development Award from the Jacobs School of Engineering at UC San Diego. The material in this paper was partially presented at the 6th Annual Learning for Dynamics & Control Conference, University of Oxford, 15-17 July, 2024 [2].
1 Introduction
There has been a surging interest in utilizing data-driven techniques to control systems with unknown dynamics [3, 4, 5, 6]. Existing data-driven methods can be generally categorized into indirect and direct control techniques. The indirect data-driven control approaches typically include the sequential system identification (system ID) and model-based control [7, 8, 9, 10]. This two-stage control pipeline has been widely used, especially for linear systems. More recently, the Koopman operator has been leveraged to construct models of unknown nonlinear systems [11, 12, 13], but the accuracy of such models highly depends on the choice of lifting functions that are non-trivial to select [14]. In contrast, direct data-driven control methods bypass system identification altogether and design control strategies directly from input-output data, offering practitioners a potentially more attractive end-to-end alternative [4, 1, 6].
One popular direct approach is the data-driven predictive control (DDPC) [4], which utilizes Willem’s fundamental lemma [15] to construct a data-driven representation of the system and use it in receding horizon control. One of the earliest formulations is the so-called DeePC, proposed in [16]. DeePC is initially established for deterministic linear time-invariant (LTI) systems, and its equivalence with subspace predictive control (SPC) has been discussed in [17]. Subsequent works [18, 19] have further investigated terminal constraint design for the closed-loop stability in LTI systems.
DeePC and its general DDPC variants have demonstrated promising experimental results for controlling systems beyond LTI settings [20, 21, 22, 23, 24]. The recent work [25] has established some theoretical guarantees for DDPC in nonlinear systems. For non-deterministic or nonlinear systems, both offline and online collected data are needed to increase the width (i.e., column number) and depth (i.e., row number) of the trajectory library, so that an accurate enough data-driven representation can be constructed. The benefits of increasing its width are well-recognized in the literature [26, 21, 22, 23], and the recent works have further emphasized the importance of enlarging its depth [14, 27]. Moreover, appropriate regularizations are very important for DDPC to ensure reliable closed-loop performance [28, 1, 4].
Figure 1: Schematic of data-driven predictive control (DDPC), which starts by collecting data (usually noisy) from the real system. Indirect methods identify a parametric model, while DDPC forms a Hankel matrix as the trajectory library for predictive control. Our bi-level formulation (7) integrates system ID techniques for trajectory library in DDPC. We introduce a series of convex relaxation (23), (29), (33) and approximation (38) for the bi-level formulation.
Since the introduction of DeePC in [16], numerous regularization strategies for DDPC have been developed [28, 1, 29, 30, 21]. Early approaches based on $\ell_{1}$- and $\ell_{2}$-norm penalties were primarily heuristic, which aim to improve empirical control performance and ensure numerical stability for systems beyond LTI settings [16, 29, 30, 21]. A novel projection-based regularizer was later proposed in [1], establishing a formal connection between DeePC and SPC; in fact, the two formulations are equivalent when the weighting parameter is sufficiently large. Another line of work introduced $\gamma$-DDPC [28], which reformulates the original data-driven representation via an LQ factorization, introducing a new decision variable $\gamma$. Regularization is then applied directly to $\gamma$, which offers a potentially transparent interpretation of the regularization effects. Moreover, it is shown in [31, Theorem 2] that $\gamma$-DDPC with $\ell_{2}$ regularization and a sufficiently large weight is equivalent to SPC. This framework was further extended in [31] by incorporating causality constraints and designing an associated regularizer. More recently, [32] proposed a maximum-likelihood estimator that characterizes future input-output trajectories through an iterative algorithm, thereby eliminating the need for manual parameter selection.
The advantages and limitations of direct and indirect data-driven methods have been extensively discussed; see the editorial column [6] for an excellent overview. Recent works have further examined the role of regularization in connecting these two paradigms. A notable contribution is [1], which analyzes the effect of regularization in a principled way via a bi-level optimization framework. In this setting, indirect data-driven control is formulated as a bi-level problem involving both identification and control, and many regularized variants of DDPC (e.g., $\ell_{1}$-norm penalties and projection-based regularizers) can be interpreted as convex relaxations of this formulation. A similar analysis has been carried out for $\gamma$-DDPC in [31], where the identification task serves as the inner problem. An alternative perspective is established in [33], which interprets the regularization term as an implicit predictor: the regularizer implicitly selects a model class for the data-driven representation and reduces model complexity. The recent work [34] introduces the concept of the final control error (FCE), defined as the expected control cost with respect to the model distribution given pre-collected data. The proposed FCE-DDPC minimizes this measure and includes certain regularized DDPC and $\gamma$-DDPC as suboptimal instances [34].
In this paper, we adopt a bi-level optimization framework, similar to [1], to study the interplay between direct and indirect data-driven control. In this formulation, data pre-processing is modeled as an inner optimization problem (identification), while predictive control is formulated as an outer optimization problem (online control). Figure 1 illustrates the overall process. Our main technical contributions are as follows.
First, we highlight the role of the Hankel trajectory library $H$ as a non-parametric model within the bi-level framework. This choice not only facilitates the incorporation of common system ID constraints (e.g., SPC [35], low-rank approximation [36], and causal models [37]) into the inner problem, but also enables the use of penalty methods to reduce the bi-level formulation to a single-level problem (Theorem 1). Second, under this bi-level framework, we establish three regularized DDPC formulations as convex relaxations of their corresponding bi-level optimization problems (Theorems 2, 3, 4). In each case, the explicit projection of data onto LTI identification constraints is replaced with suitable regularizations that account for implicit identification. While this perspective aligns with [1], we refine the analysis by clarifying conceptual inaccuracies, providing elementary self-contained proofs, and introducing a new causality-based regularizer (Theorem 3). Finally, all three regularized DDPC formulations only implicitly enforce a subset of system ID constraints while neglecting others. To address this limitation, we propose an optimality-based variant of DDPC, O-DDPC, which approximately solves the inner identification problem with all system ID constraints via an iterative algorithm. This leads to a refined data-driven system representation that captures linearity, causality, and the dominant trajectory space. Numerical experiments demonstrate that O-DDPC outperforms existing regularized DDPC approaches in reducing both bias and variance errors.
The rest of this paper is structured as follows. Section 2 reviews the preliminaries and introduces the problem statement. Section 3 discusses the use of the trajectory matrix as a predictive model in both indirect and direct approaches. In Section 4, we present three regularized DDPC formulations as convex relaxations of suitable bi-level problems. Section 5 develops the proposed optimality-based DDPC. Numerical results are reported in Section 6, and Section 7 concludes the paper. Some technical proofs are provided in the appendix.
2 Preliminaries
This section reviews model-based predictive control, the fundamental Lemma [15] from a behavioral perspective, and a basic data-driven predictive control in [16].
2.1 LTI systems and model predictive control
Consider a discrete-time LTI system:
$x(k+1)$ $=Ax(k)+Bu(k),$ (1)
$y(k)$ $=Cx(k)+Du(k),$
where the state, input, output at time $k$ are $x(k)\in\mathbb{R}^{n}$, $u(k)\in\mathbb{R}^{m}$, and $y(k)\in\mathbb{R}^{p}$, respectively. Throughout this paper, we assume that $(A,B)$ is controllable and $(C,A)$ is observable. The lag of the LTI system (1) is defined as the minimum integer $l\in\mathbb{N}$ such that its observability matrix $\textnormal{col}(C,CA,\ldots,CA^{l-1})$ has full column rank $n$. It is known that $l\leq n$ when $(C,A)$ is observable.
Given a desired reference trajectory $y_{\textrm{r}}\in\mathbb{R}^{pN}$ with horizon $N>0$, input constraint set $\mathcal{U}\subseteq\mathbb{R}^{m}$, output constraint set $\mathcal{Y}\subseteq\mathbb{R}^{p}$, we aim to design control inputs such that the system output tracks the reference trajectory. In model predictive control (MPC), this is achieved, at time $t$, by solving the receding horizon predictive control
$\min_{\bar{x},\bar{u},\bar{y}}\;\,$ $\sum_{k=t}^{t+N-1}\big{(}\|\bar{y}(k)-y_{\textrm{r}}(k)\|_{Q}^{2}+\|\bar{u}(k)\|_{R}^{2}\big{)}$
$\mathrm{subject~to}\;\,$ $\bar{x}(k+1)=A\bar{x}(k)+B\bar{u}(k),$ (2a)
$\bar{y}(k)={C}\bar{x}(k)+{D}\bar{u}(k),$ (2b)
$\bar{x}(t)=x_{t},$ (2c)
$\bar{u}(k)\!\in\!\mathcal{U},\bar{y}(k)\!\in\!\mathcal{Y},k=t,\ldots,t\!+\!N\!-\!1,$ (2d)
where $x_{t}\!\in\!\mathbb{R}^{n}$ is the initial state at time $t$, $\bar{x}(k),\bar{u}(k)$ and $\bar{y}(k)$ denote the predicted state, input, and output at time $k$, and $R$ and $Q$ are positive definite cost matrices. According to (2a)-(2c), $\{\bar{x}(k),\bar{u}(k),\bar{y}(k)\}_{k=t}^{t+N-1}$ is a predicted trajectory of the system (1) with an initial state $x_{t}$ at time $t$ using the system model (1). Equation (2d) enforces the constraints on the predicted trajectory, and we assume $\mathcal{U}$ and $\mathcal{Y}$ are convex sets. To simplify notation and without loss of generality, we consider a regulation problem (i.e., $y_{\textrm{r}}=\mathbb{0}_{pN}$) for the rest of the discussions.
As commonly used in MPC, the first optimal input of (2) is applied to the system (1) and a new optimal input is computed based on a new state measurement at the next step $t+1$. Guarantees on closed-loop performance (e.g., closed-loop stability and constraint satisfaction) can be ensured either by a sufficiently long horizon $L$ or by incorporating suitable terminal ingredients [38]. It is clear that (2) is a convex optimization problem (it is a quadratic program when $\mathcal{U}$ and $\mathcal{Y}$ are polytope), which admits an efficient solution when the model for (1) is known, i.e., matrices $A$, $B$, $C$ and $D$ are known.
In this work, we focus on the case where both the system model (1) and the initial condition $x_{t}$ are unknown. Instead, we have access to
  1. 1.
    offline data, i.e., a length-$T$ pre-collected input and output trajectory $u_{\textnormal{d}}=\textnormal{col}(u_{\textnormal{d}}(1),\ldots,u_{\textnormal{d}}(T))\in\mathbb{R}^{mT}$, $y_{\textnormal{d}}=\textnormal{col}(y_{\textnormal{d}}(1)$, $\ldots,y_{\textnormal{d}}(T))\in\mathbb{R}^{pT}$ from (1);
  2. 2.
    online data, i.e., the most recent past input and output sequence of length-$T_{\textnormal{ini}}$.
Then, problem (2) can be implemented by either indirect system identification and model-based control [39, 38] or the recent direct data-driven predictive control, such as the so-called DeePC [16] and its related approaches [1, 4, 28]. The advantages and limitations of these two classes of methods have been extensively discussed in the literature; see, for example, the recent editorial column in [6].
2.2 Direct data-driven predictive control
We here review the notion of persistent excitation (PE) to ensure the offline data is sufficiently rich.
Definition 1 (Persistently Exciting).
A sequence of data points $\omega=\textnormal{col}(\omega(1),\omega(2),\ldots,\omega(T))$ of the length $T$ is persistently exciting (PE) of order $L$ ($L<T$) if its associated Hankel matrix with depth $L$,
[…]
has full row rank.
The following result from [15], commonly known as the Fundamental lemma, forms the foundation of many recent results of direct data-driven predictive control.
Lemma 1 (Fundamental Lemma [15]).
Suppose (1) is controllable. Given a length-$T$ input/output trajectory $u_{\textnormal{d}}\in\mathbb{R}^{mT}$ and $y_{\textnormal{d}}\in\mathbb{R}^{pT}$ where $u_{\textnormal{d}}$ is PE of order $L+n$, then a length-$L$ input/output sequence $\{u_{\textnormal{s}}(k),y_{\textnormal{s}}(k)\}_{k=0}^{L-1}$ is a valid trajectory of (1) if and only if there exists a $g\in\mathbb{R}^{T-L+1}$ such that
[…]
If $L$ is not smaller than the lag of the system (1), matrix $\textnormal{col}(\mathcal{H}_{L}(u_{\textnormal{d}}),\mathcal{H}_{L}(y_{\textnormal{d}}))$ has rank $mL+n$.
The fundamental lemma gives a parameterization of all finite-dimensional input/output trajectories of (1) using only one offline trajectory $\{u_{\textnormal{d}},y_{\textnormal{d}}\}$. In particular, the image of the Hankel matrix in (3) is the same as the set of all system trajectories of length $L$. This serves as the foundation of many recent direct data-driven analysis and control methods; see [4, 40, 41] for excellent surveys.
With the fundamental lemma, we can use (3) to build a predictive model to replace (2a)-(2c). This is the basic idea in DeePC [16]. In particular, the Hankel matrix formed by the offline data in (3) is partitioned as
$\begin{bmatrix}U_{\textrm{P}}\\ U_{\textrm{F}}\end{bmatrix}:=\mathcal{H}_{L}(u_{\textrm{d}}),\quad\begin{bmatrix}Y_{\textrm{P}}\\ Y_{\textrm{F}}\end{bmatrix}:=\mathcal{H}_{L}(y_{\textrm{d}}),$ (4)
where $U_{\textnormal{P}}$ and $U_{\textnormal{F}}$ consist the first $T_{\textnormal{ini}}$ rows and the last $N$ rows of $\mathcal{H}_{L}(u_{\textnormal{d}})$, respectively (similarly for $Y_{\textnormal{P}}$ and $Y_{\textnormal{F}}$). The length of offline data satisfies $L\!=\!T_{\textnormal{ini}}\!+\!N$. We denote the most recent past input trajectory of length $T_{\textnormal{ini}}$ and the future input trajectory of length $N$, respectively, as
$u_{\textnormal{ini}}$ $=\textnormal{col}(u(t-T_{\textrm{ini}}),u(t-T_{\textrm{ini}}+1),\ldots,u(t-1)),$
$\bar{u}$ $=\textnormal{col}(\bar{u}(t),\bar{u}(t+1),\ldots,\bar{u}(t+N-1)),$
(similarly for $y_{\textrm{ini}},\bar{y}$). Then, Lemma 1 ensures that the sequence $\textnormal{col}(u_{\textrm{ini}},y_{\textrm{ini}},\bar{u},\bar{y})$ is a valid trajectory of (1) if and only if there exists $g\in\mathbb{R}^{T-T_{\textrm{ini}}-N+1}$ such that
$H_{\mathrm{d}}g=\textnormal{col}(u_{\textrm{ini}},y_{\textrm{ini}},\bar{u},\bar{y}),$ (5a)
where, for notational simplicity, we denote
$H_{\mathrm{d}}:=\textnormal{col}(U_{\textnormal{P}},Y_{\textnormal{P}},U_{\textnormal{F}},Y_{\textnormal{F}}),\vskip-5.69054pt$ (5b)
as the Hankel matrix associated with pre-collected data $\{u_{\mathrm{d}},y_{\mathrm{d}}\}$; see (4).
If $T_{\mathrm{ini}}$ is larger or equal to the lag of (1), $\bar{y}$ is unique given any $u_{\textrm{ini}}$, $y_{\textrm{ini}}$ and $\bar{u}$ in (5).
The basic DeePC [16] utilizes (5) as the data-driven representation of (2a)-(2c) and reformulate problem (2) as
$\min_{g,\bar{u},\bar{y}}$ $\sum_{k=t}^{t+N-1}\big{(}\|\bar{y}(k)\|_{Q}^{2}+\|\bar{u}(k)\|_{R}^{2}\big{)}$ (6)
[…]
where we slightly abuse the notation and use $\bar{u}\in\mathcal{U},\bar{y}\in\mathcal{Y}$ to denote input/output constraints (2d).
2.3 Direct vs. indirect data-driven control
For LTI systems with noise-free data, model-based control (2) and DeePC (6) are equivalent (see [16, Theorem 5.1]), thanks to the fundamental lemma. In this case, the Hankel matrix $H_{\mathrm{d}}$ in (5b), also referred to as the trajectory matrix (since each of its columns represents a valid system trajectory), serves as a non-parametric model.
However, data collected from practical systems is rarely noise-free. In particular, the offline data $u_{\mathrm{d}},y_{\mathrm{d}}$ and the resulting trajectory library $H_{\mathrm{d}}$ in (5b) may be corrupted by 1) “variance” noises that enter the process dynamics and output measurement, 2) and “bias” errors induced by nonlinear dynamics beyond LTI [1]. To address these issues, various regularization and data preprocessing techniques have been proposed to extend the basic DeePC (6). These include $l_{1}$$l_{2}$ regularization [1], $\gamma$-DDPC [31], low-rank approximation [36], singular-value decomposition [26]. While some works [26, 1, 42, 31] have explored the relationship among different schemes, most of them are carried out on a case-by-case basis. One notable exception is [1], which employs a principled bi-level optimization framework to investigate the interplay between direct and indirect data-driven control.
We here adopt the same bi-level optimization principle as [1] to systematically explore the effects of regularization in direct Data-driven Predictive Control (DDPC). In particular, we consider a nested bi-level DDPC,
$\min_{\begin{subarray}{c}g,\sigma_{y}\in\Gamma,\\ u\in\mathcal{U},y\in\mathcal{Y}\end{subarray}}\quad$ $\|y\|_{Q}^{2}+\|u\|_{R}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2},$
$\mathrm{subject~to}\quad$ ${H}^{\star}g=\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y},u,y),$ (7a)
$\textrm{where}\quad{H}^{\star}\in\arg\min_{{H}\in\mathcal{S}}\;J_{\mathrm{id}}({H},H_{\mathrm{d}}),$ (7b)
where we process the offline data $H_{\mathrm{d}}$ (see (5)) in the inner identification (7b) before using it for the outer online control (7a). This bi-level formulation, often referred to as indirect DDPC, is modular and consists of two well-separated components: an inner identification layer and an outer model-based predictive control layer. Before discussing its connections and distinctions with existing schemes, let us further clarify the notation in (7). In the outer problem, the term $\|y\|_{Q}^{2}+\|u\|_{R}^{2}$ denotes the usual one $\sum_{k=t}^{t+N-1}\big{(}\|y(k)\|_{Q}^{2}+\|u(k)\|_{R}^{2}\big{)}$, and we have introduced a slack variable $\sigma_{y}$ with constraint $\Gamma$ and a regularization term $\|\sigma_{y}\|_{2}^{2}$ with $\lambda_{y}>0$. This ensures constraint feasibility despite noise, as used in [4, 18]. In the inner problem, $J_{\mathrm{id}}(\cdot,\cdot)$ denotes a suitable identification loss, and the constraint ${H}\in\mathcal{S}$ enforces some prior data structures from LTI systems. We will clarify our choice of ${H}\in\mathcal{S}$ in Section 3.
Remark 1 (Comparison with [1]).
The bi-level DDPC (7) is inspired by the recent work [1], which argues that “direct and regularized data-driven control can be derived as a convex relaxation of the indirect approach”. While adopting the same general framework, our work introduces two key distinctions. Conceptually, the bi-level DDPC (7) explicitly emphasizes the use of a non-parametric trajectory matrix as the predictive model, whereas [1] primarily considers a more general and abstract behavioral LTI setting in the inner problem. Technically, the analysis in [1] relies on arguments that require additional assumptions or clarifications, and some claims may not hold without these refinements. Also, the exact penalty arguments in [1] rely on [43, Proposition 2.4.3], which requires adaptation in the DDPC context. Inspired by the concept of partial calmness, our proof is elementary, transparent, and self-contained. We will clarify these points in more detail in Sections 3 and 4.
3 Trajectory matrix as a predictive model: indirect and direct approaches
In this section, we first detail the inner identification (7b) and then discuss how to bridge the bi-level DDPC (7) with a single-level optimization via penalty methods.
3.1 Behavior-based identification for predictive control
From the offline noisy data $H_{\mathrm{d}}$, we aim to get a new trajectory library $H^{\star}$ for predictive control
${H}^{\star}:=\textnormal{col}({U}_{\textnormal{P}}^{\star},{Y}_{\textnormal{P}}^{\star},{U}_{\textnormal{F}}^{\star},{Y}_{\textnormal{F}}^{\star}),$
where each column of $H^{\star}$ is a “purified” trajectory of the system. Let $\mathcal{B}_{L}$ (where $L=T_{\textnormal{ini}}+N$) denotes the space of all possible length-$L$ trajectories of an LTI system, i.e.,
[…]
Ideally, we may consider the following problem
$H^{\star}=\operatorname*{argmin}_{H}$ $J_{\mathrm{id}}({H},H_{\mathrm{d}})=\|H-H_{\mathrm{d}}\|_{F}^{2}$ (8)
$\mathrm{subject~to}$ $\mathrm{Im}\,{H}\subseteq\mathcal{B}_{L}.$
If the offline data $H_{\mathrm{d}}$ is noise-free and satisfies the persistent excitation, the fundamental lemma ensures that $\mathrm{Im}\,{H}_{\mathrm{d}}=\mathcal{B}_{L}$. In this case, the optimal solution to (8) is $H^{\star}={H}_{\mathrm{d}}$. In general, the constraint $\mathrm{Im}\,{H}\subseteq\mathcal{B}_{L}$ is too abstract and not tractable. For an explicit expression, we have several necessary conditions.
  • First, the trajectory matrix $H$ should be of low rank, i.e., we have $\mathrm{rank}\,H^{\star}\leq mL+n$.
  • Second, the trajectory matrix $H$ should satisfy linearity, i.e., the future output is a linear function of the past data and future input, and we have
    ${Y}_{\textnormal{F}}^{\star}=K\ \textnormal{col}(U_{\textnormal{P}},Y_{\textnormal{P}},U_{\textnormal{F}}),$ (9)
    for some coefficient matrix $K$.
  • Third, the trajectory matrix $H$ should satisfy causality, i.e., the coefficient matrix $K$ in (9) should have a block partition $K=\begin{bmatrix}K_{\mathrm{p}}\;K_{\mathrm{f}}\end{bmatrix},\ K_{\mathrm{f}}\in\mathcal{L}$, where $K_{\mathrm{f}}\in\mathcal{L}$ encodes a block-low triangle structure for causality.
  • Finally, we may also impose a Hankel structure on $H^{\star}\in\mathcal{H}$ if each column is a one-step shifted trajectory.
We then relax (8) with the following problem
$({H}^{\star},K^{\star})\in\operatorname*{argmin}_{\tilde{H},K}\quad$ $\|\tilde{H}-H_{\mathrm{d}}\|_{F}^{2}$
$\mathrm{subject~to}\quad$ $\mathrm{rank}(\tilde{H})=mL+n,$ (10a)
$\tilde{Y}_{\textnormal{F}}=K\ \textnormal{col}(U_{\textnormal{P}},Y_{\textnormal{P}},U_{\textnormal{F}}),$ (10b)
$K=\begin{bmatrix}K_{\mathrm{p}}\quad K_{\mathrm{f}}\end{bmatrix},\ K_{\mathrm{f}}\in\mathcal{L},$ (10c)
$\tilde{H}\in\mathcal{H}$. (10d)
We thus consider (10) as our inner identification in (7b). Note that the requirements of linearity (10b) and causality (10c) are typically used in causal SPC [35, 17], and the requirement of low rank (10a) and Hankel structure (10d) are often employed in low-rank approximation [36]. When the offline data $H_{\mathrm{d}}$ in (5) is collected from an LTI system (1) with noise-free data, the unique optimal solution to (10) is trivially ${H}^{*}=H_{\mathrm{d}}$. If $H_{\mathrm{d}}$ contains “variance” noises and/or “biases” errors, the identification problem (10) becomes challenging to solve.
Remark 2.
[…]
3.2 A direct approach via penalty methods
With the inner identification (10), the bi-level formulation (7) is an indirect DDPC. Using penalty methods, the bi-level structure can be transformed into a single-level problem. This is a standard idea in bi-level optimization [44] and has been recently used to analyze direct data-driven control in [1]. We adopt a similar strategy here, but also highlight a subtle point that leads to some conceptual inaccuracies in [1].
A key step involves replacing the inner optimization problem (10) with equivalent (but implicit) constraints, which are then incorporated into the objective function via penalty methods. One conceptually simple way is to assume the existence of a set of optimality conditions $H\in\mathcal{C}_{\mathrm{opt}},$ where $\mathcal{C}_{\mathrm{opt}}$ denotes the optimality constraints for the inner problem (10). Then, we can write problem (7) equivalently as
$\min_{\begin{subarray}{c}g,\sigma_{y}\in\Gamma,{H},\\ u\in\mathcal{U},y\in\mathcal{Y}\end{subarray}}\quad$ $\|y\|_{Q}^{2}+\|u\|_{R}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2}$
$\mathrm{subject~to}\quad$ ${H}g=\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y},u,y),$ (11a)
${H}\in\mathcal{C}_{\mathrm{opt}}.$ (11b)
We note that (11) is only of a conceptual use at this stage, as (11b) may not be known explicitly. We next consider a continuous penalty function satisfying
$\left\{\begin{aligned} p(H)=0,&\quad\textrm{if}\ {H}\in\mathcal{C}_{\mathrm{opt}},\\ p(H)>0,&\quad\textrm{if}\ {H}\notin\mathcal{C}_{\mathrm{opt}}.\end{aligned}\right$. (12)
Now, we can write a penalized problem
$\min_{\begin{subarray}{c}g,\sigma_{y}\in\Gamma,{H},\\ u\in\mathcal{U},y\in\mathcal{Y}\end{subarray}}$ $\|y\|_{Q}^{2}+\|u\|_{R}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2}+\lambda p(H)$ (13)
$\mathrm{subject~to}$ ${H}g=\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y},u,y).$
Under a mild assumption (see e.g., [45, Theorem 6.6]), we can ensure that as the penalty parameter $\lambda\to\infty$, the optimal solution of (13) converges to an optimal solution of (11), i.e., the original bi-level problem (7). If we use a small penalty parameter $\lambda$, the single-level formulation (13) becomes a relaxation of (7), i.e., the optimal value of (13) is smaller than that in (7).
The reasoning above critically relies on the conceptual optimality constraint $H\in\mathcal{C}_{\mathrm{opt}}$, associated with the inner problem (7b). Dropping any constraint in (7b) results in a different optimality set, meaning that the corresponding penalized single-level problem may no longer offer a valid relaxation. This subtle but important point may be easily overlooked; for example, some relaxation statements in [1] are not fully accurate without further clarification. We consider a simple example below.
Example 1.
Consider a simple bi-level formulation
$\min_{x\in\mathbb{R}}\quad$ $x^{2}$
$\mathrm{subject~to}\quad$ $y^{*}x=1,$ (14a)
$\mathrm{where}\;y^{*}\in\arg\min_{y\in\mathbb{R}}\;y^{2}$ (14b)
$\qquad\qquad\mathrm{subject~to}\;\;2\leq y\leq 4.$ (14c)
This inner problem has an optimality condition as $y^{*}\in\mathcal{C}_{\mathrm{opt}}:=\{2\}$. We consider a penalty function p(y)=|y2|p(y)=|y-2|, which satisfies (12). Then, a penalized single-level optimization is
$\min_{x\in\mathbb{R},y\in\mathbb{R}}$ $x^{2}+\lambda|y-2|$ (15)
$\mathrm{subject~to}$ $yx=1.$
It can be verified that when $\lambda\geq\frac{1}{4}$ (see Figure 2 for illustration), (15) has the same optimal solution as (14). If $\lambda<\frac{1}{4}$, the optimal value of (15) is a lower bound of (14). If we relax the constraint (14c) as $y\leq 4$, then the inner problem has a different optimal solution $y^{*}=0$. In this case, the outer problem in (14a) becomes infeasible. This illustrates that dropping constraints from the inner problem does not necessarily reduce the optimal value of the outer problem. This nuance was overlooked in [1].
3.3 Shifting regularization from $H$ to $g$
The direct data-driven method (13) is conceptually useful but not practically implementable for two reasons: 1) the regularization is imposed on the Hankel matrix $H$, which is implicit; and 2) the constraint in (13) is bilinear in $H$ and $g$, which is computationally intractable. We here discuss how to shift the regularization from $H$ to $g$.
Figure 2: The influence of the penalty parameter $\lambda$ in (15), where we have replaced $y$ with $1/x$.
Our general idea is to select some identification constraints from (10a)-(10d) such that the inner identification problem admits a simpler and potentially explicit optimality condition $\hat{\mathcal{C}}_{\mathrm{opt}}$. Then, we can impose some constraints $h:\mathbb{R}^{\bar{n}}\rightarrow\mathbb{R}$ on $g$ such that the set
$\{(g,\sigma_{y},u,y)\mid Hg=\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y},u,y),{H}\in\hat{\mathcal{C}}_{\mathrm{opt}}\}$
is the same as
$\{(g,\sigma_{y},u,y)\mid Dg=\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y},u,y),h(g)=0\},$
where $D$ is a fixed trajectory matrix. We next move the constraint on $g$ to the cost via regularization, leading to a direct DDPC
$\min_{\begin{subarray}{c}g,\sigma_{y}\in\Gamma,\\ u\in\mathcal{U},y\in\mathcal{Y}\end{subarray}}$ $\|y\|_{Q}^{2}+\|u\|_{R}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2}+\lambda_{w}|h(g)|$ (16)
$\mathrm{subject~to}$ $Dg=\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y},u,y).$
Section 4 will detail three ways to derive an explicit form (16), including regularization using projection-based norm, causality-based norm, and $l_{1}$ norm.
Similar to (13), one may require the regularization parameter $\lambda_{w}$ in (16) to approach infinity. Fortunately, in our context, we can establish an exact penalty with a finite regularization parameter $\lambda_{w}$. In particular, (16) is a quadratic optimization problem of the form
$\min_{x_{1}\in\mathcal{X},x_{2}}$ $x_{1}^{{\mathsf{T}}}Mx_{1}$ (17)
$\mathrm{subject~to}$ $A_{1}x_{1}+A_{2}x_{2}=b,$
$\|Dx_{2}\|_{p}=0,$
where $M$ is positive semidefinite, $A_{1}$, $A_{2}$, $b$, and $D$ are problem data of compatible dimensions, $\mathcal{X}$ is a convex set, and $\|\cdot\|_{p}$ is any $p$-norm. We note that $x_{2}$ plays the role of $g$ in (16). We then consider a regularized version
$\min_{x_{1}\in\mathcal{X},x_{2}}$ $x_{1}^{{\mathsf{T}}}Mx_{1}+\lambda_{w}\|Dx_{2}\|_{p}$ (18)
$\mathrm{subject~to}$ $A_{1}x_{1}+A_{2}x_{2}=b.$
We next show that under mild assumptions, there exists a finite $\lambda_{w}>0$ such that (17) and (18) are equivalent.
Theorem 1.
Consider (17) and (18), where $\mathcal{X}$ is a convex set, $M$ is positive semidefinite, and $x_{1}^{{\mathsf{T}}}Mx_{1}$ is Lipschitz continuous with Lipschitz constant $L$ with respect to $\|\cdot\|_{p}$ over $\mathcal{X}$. If there exists an optimal solution $(x_{1}^{*},x_{2}^{*})$ for (17) such that $x_{1}^{*}$ is an interior point of $\mathcal{X}$, then there exists $\lambda_{w}^{*}>0$, such that (17) and (18) are equivalent for all $\lambda_{w}>\lambda_{w}^{*}$, i.e., they have same optimal solutions.
One key fact in this theorem is that we can find a finite penalty parameter such that (17) and (18) are equivalent. Theorem 1 confirms that $\|Dx_{2}\|_{p}$ is an exact penalty function for the quadratic problem (17). To our knowledge, this result is new and might be of independent interest. The key proof idea is to show that the optimal solution of (17) is also a local minimizer of (18). Since (18) is a convex problem, any local minimizer is also a global minimizer. Consequently, $\lambda_{w}\|Dx\|_{p}$ serves as an exact penalty term for some finite weight $\lambda_{w}>\lambda_{w}^{\star}$. The proof details are technically involved, and we present the full proof in Appendix A.
To illustrate Theorem 1, we randomly generate problem instances of (17) and (18). As shown in Figure 3, all $p$-norms render $\|Dx\|_{p}$ an exact penalty. However, smaller values of $p$ lead to faster convergence of the optimal value of (18) toward that of (17) as $\lambda_{w}$ increases. Since $\|Dx\|_{p}$ decreases monotonically with $p$, larger $p$ values impose weaker penalties and therefore require larger $\lambda_{w}$ to guarantee equivalence between the two problems.
4 Regularization in direct data-driven control via convex relaxations
In this section, we apply Theorem 1 and derive three direct DDPC strategies via convex relaxations, including regularization using projection-based norm, causality-based norm, and $l_{1}$ norm. The projection-based norm and $l_{1}$ norm are widely used in the literature (see [1] and its references), and the causality-based norm is new.
As indicated in Section 3.3, our general idea is to select some identification constraints from (10a)-(10d) such that the inner identification problem admits a simpler and potentially explicit optimality condition.
4.1 Projection-based norm and a Hankel-form SPC
We first consider the following bi-level problem
$\min_{\begin{subarray}{c}g,\sigma_{y}\in\Gamma,\\ u\in\mathcal{U},y\in\mathcal{Y}\end{subarray}}\quad$ $\|y\|_{Q}^{2}+\|u\|_{R}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2}$
$\mathrm{subject~to}\quad$ ${H}^{*}g=\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y},u,y),$ (19)
$\textrm{where}\;\;{H}^{*}\in\operatorname*{argmin}_{\tilde{H},K}\;\;\|\tilde{H}-H_{\mathrm{d}}\|_{F}^{2}$
[…]
which only considers the SPC constraint (10b) for the inner identification. This bi-level DDPC (19) is much simpler than the general one (7). However, similar to Example 1, the optimal value of (19) may not be lower than that of (7); in other words, (19) is not a relaxation of (7).
Figure 3: Numerical illustration of Theorem 1. The optimal value of (18) increases with larger weight $\lambda_{w}$ across different $p$-norms ($p=1,2,\infty$). In all cases, once $\lambda_{w}$ exceeds a threshold, the optimal value of (18) coincides with that of (17).
It is not difficult to derive the unique optimal solution $H^{*}_{\textnormal{s}}$ for the inner problem in (19), which can be obtained via Moore–Penrose inverse as ${H}^{*}_{\textnormal{s}}=\textnormal{col}(U_{\textnormal{P}},Y_{\textnormal{P}},U_{\textnormal{F}},M)$ where $M=Y_{\textnormal{F}}\Pi_{1}$ with
$\Pi_{1}:=H_{1}^{\dagger}H_{1},\qquad H_{1}:=\textnormal{col}(U_{\textnormal{P}},Y_{\textnormal{P}},U_{\textnormal{F}}).$ (20)
Then, the bi-level problem (19) is equivalent to
$\min_{\begin{subarray}{c}g,\sigma_{y}\in\Gamma,\\ u\in\mathcal{U},y\in\mathcal{Y}\end{subarray}}$ $\|u\|_{R}^{2}+\|y\|_{Q}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2}$ (21)
$\mathrm{subject~to}$ ${H}^{*}_{\textnormal{s}}g=\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y},u,y).$
Note that (21) differs slightly from the classical SPC formulation and we call it Hankel-form SPC. In particular, the classical SPC does not involve the use of a Hankel matrix; instead, it employs an explicit multi-step predictor (also see Remark 2). For self-completeness, we provide further details in Appendix B.1.
Following the discussion in Section 3.3, we here fix $H=H_{\mathrm{d}}$ and impose a constraint on $g$ to get an equivalent reformulation for (21). In this case, we use a projection-based norm $\|(I-\Pi_{1})g\|_{2}$ as in [1], and consider
$\min_{\begin{subarray}{c}g,\sigma_{y}\in\Gamma,\\ u\in\mathcal{U},y\in\mathcal{Y}\end{subarray}}\quad$ $\|y\|_{Q}^{2}+\|u\|_{R}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2}$
$\mathrm{subject~to}\quad$ $H_{\mathrm{d}}g=\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y},u,y),$ (22a)
$\|(I-\Pi_{1})g\|_{2}=0,$ (22b)
where the matrix $\Pi_{1}$ is defined in (20).
Proposition 1.
Suppose $Q\succ 0,R\succ 0,\lambda_{y}>0$ and that $\Gamma,\mathcal{U},$ and $\mathcal{Y}$ are convex sets. Fix any data matrix $H_{\textnormal{d}}$ in (5b). Then, the bi-level problem (19) has the same optimal value as the single-level problems (21) and (22). Moreover, all three problems admit the same unique optimal solution $u^{*},y^{*}$ and $\sigma_{y}^{*}$.
The proof is provided in Appendix B.2. The key idea is to show that feasible regions for decision variables $u,y,\sigma_{y}$ are the same for (19), (21) and (22). Note that the data matrix $H_{\textnormal{d}}$ in Proposition 1 can be arbitrary and may come from a nonlinear system.
It is now clear that (22) is in the form of (17). Thus, applying Theorem 1 directly leads to the following result.
Corollary 1.
Suppose $Q\succ 0,R\succ 0,\lambda_{y}>0$ and $\Gamma,\mathcal{U},$ and $\mathcal{Y}$ are compact and convex sets. If the optimal solution $(\sigma_{y}^{\star},u^{\star},y^{\star})$ to (22) is an interior point of $\Gamma\times\mathcal{U}\times\mathcal{Y}$, then there exists $\lambda_{g}^{*}>0$ such that
$\min_{\begin{subarray}{c}g,\sigma_{y}\in\Gamma,\\ u\in\mathcal{U},y\in\mathcal{Y}\end{subarray}}$ $\|u\|_{R}^{2}+\|y\|_{Q}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2}$ (23)
$+\lambda_{g}\|(I-\Pi_{1})g\|_{2}$
$\mathrm{subject~to}$ $H_{\textnormal{d}}g=\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y},u,y),$
is equivalent to (22) for all $\lambda_{g}>\lambda_{g}^{*}$. In other words, they have the same optimal cost value and optimal solution.
We provide some details in Appendix B.4. We call (23) as Linearity-based DDPC (L-DDPC). Combining Corollary 1 and Proposition 1 gives us the following result.
Theorem 2.
Consider the bi-level DDPC (19) and the single-level L-DDPC (23). Suppose $Q\succ 0,R\succ 0,\lambda_{y}>0$ and $\Gamma,\mathcal{U},\mathcal{Y}$ are compact and convex sets and the optimal solution $(\sigma_{y}^{*},u^{*},y^{*})$ for (19) is an interior point of $\Gamma\times\mathcal{U}\times\mathcal{Y}$. Then, for any nonnegative $\lambda_{g}\geq 0$, (23) is a convex relaxation of (19), that is:
  1. 1.
    Problem (23) is convex;
  2. 2.
    Any feasible $\sigma_{y},u,y$ for (19) is feasible to (23);
  3. 3.
    The optimal value of (23) is smaller than or equal to that of (19).
Our Theorem 2 is closely related to [1, Theorem 4.6], but there are two key differences. Conceptually, as illustrated in Example 1, the L-DDPC (23) is only a convex relaxation of (19), and not necessarily of the original bi-level problem (7). This subtle distinction was overlooked in [1, Theorem 4.6]. Technically, we establish that (23) serves as a convex relaxation of (19) for any nonnegative parameter $\lambda_{g}\geq 0$, rather than only for sufficiently small $\lambda_{g}$ as in [1, Theorem 4.6]. The threshold $\lambda_{g}^{*}$ that makes (19) and (23) equivalent depends on both the Lipschitz constant of the objective function and problem data (e.g., $H_{\textnormal{d}}$) rather than only on the Lipschitz constant as in [1, Theorem 4.6]. Moreover, our proof is elementary, relying on Theorem 1 and Proposition 1, which are inspired by the theoretical condition of partial calmness. We avoid utilizing [43, Proposition 2.4.3] as in [1, Theorem4.6], whose application in the DDPC setting may require additional adaptations.
4.2 Causality-based norm and a Causal SPC
We next consider the following bi-level problem
$\min_{\begin{subarray}{c}g,\sigma_{y}\in\Gamma,\\ u\in\mathcal{U},y\in\mathcal{Y}\end{subarray}}\quad$ $\|y\|_{Q}^{2}+\|u\|_{R}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2}$
$\mathrm{subject~to}\quad$ ${H}^{*}g=\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y},u,y),$ (24)
$\textrm{where}\;\;{H}^{*}\in\operatorname*{argmin}_{\tilde{H},K}\;\;\|\tilde{H}-H_{\mathrm{d}}\|_{F}^{2}$
[…]
which considers the SPC constraint (10b) and the causality constraint (10c) for the inner identification. If $H_{\textnormal{d}}$ has full row rank, the optimal solution $H^{*}_{\textnormal{sc}}$ for the inner problem in (24) is unique [31, Lemma 2], which can be obtained via its optimal solution $K^{*}_{\textnormal{sc}}$, derived as
$K_{\textnormal{sc}}^{*}=\begin{bmatrix}L_{31}&L_{32}^{*}\end{bmatrix}\begin{bmatrix}L_{11}&\mathbb{0}\\ L_{21}&L_{22}\end{bmatrix}^{-1},$
where $L_{ij}$ comes from the LQ factorization of $H_{\textnormal{d}}$, i.e.,
$H_{\textnormal{d}}=\begin{bmatrix}L_{11}&\mathbb{0}&\mathbb{0}\\ L_{21}&L_{22}&\mathbb{0}\\ L_{31}&L_{32}&L_{33}\end{bmatrix}\begin{bmatrix}Q_{1}\\ Q_{2}\\ Q_{3}\end{bmatrix}$
and $L_{32}^{*}$ is the lower-block triangular part of $L_{32}\in\mathbb{R}^{pL\times mL}$. Then, the optimal $H_{\textnormal{sc}}^{*}$ for the inner problem in (24) can be written as
[…]
We can equivalently formulate (24) as
$\min_{\begin{subarray}{c}g,\sigma_{y}\in\Gamma,\\ u\in\mathcal{U},y\in\mathcal{Y}\end{subarray}}$ $\|u\|_{R}^{2}+\|y\|_{Q}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2}$ (26)
$\mathrm{subject~to}$ $H_{\textnormal{sc}}^{*}g=\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y},u,y).$
Following the discussion in Section 3.3, we next fix $H$ to a new trajectory library and set a suitable constraint on $g$ to obtain an equivalent reformulation for (26). For this, we denote the difference between $L_{32}$ and its lower-block triangular part $L_{32}^{*}$ as $L_{32}^{\prime}:=L_{32}-L_{32}^{*}$ and the new trajectory library $\hat{H}$ is
$\hat{H}=\begin{bmatrix}L_{11}&\mathbb{0}&\mathbb{0}&\mathbb{0}\\ L_{21}&L_{22}&\mathbb{0}&\mathbb{0}\\ L_{31}&L_{32}^{*}&L_{33}&L_{32}^{\prime}\end{bmatrix}\!\begin{bmatrix}Q_{1}\\ Q_{2}\\ Q_{3}\\ Q^{*}\end{bmatrix}\vskip-5.69054pt$ (27)
where $Q^{*}$ has orthonormal rows and $Q^{*}Q_{i}^{{\mathsf{T}}}=\mathbb{0}$ for $i=1,2,3$. We further consider the norm $\|Q_{\textnormal{c}}g\|_{2}$ where $Q_{\textnormal{c}}=\textnormal{col}(Q_{3},Q^{*})$. Then, we have
$\min_{\begin{subarray}{c}g,\sigma_{y}\in\Gamma,\\ u\in\mathcal{U},y\in\mathcal{Y}\end{subarray}}\quad$ $\|y\|_{Q}^{2}+\|u\|_{R}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2}$
$\mathrm{subject~to}\quad$ $\hat{H}g=\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y},u,y),$ (28a)
$\|Q_{\textnormal{c}}g\|_{2}=0.$ (28b)
Our next result shows that the optimal solutions of (24), (26) and (28) are all the same.
Proposition 2.
Suppose $Q\succ 0,R\succ 0,\lambda_{y}>0$ and $\Gamma,\mathcal{U},\mathcal{Y}$ are convex sets. Fix any data matrix $H_{\textnormal{d}}$ in (5b) with full row rank and sufficiently large column number. The bi-level problem (24) has the same optimal value as the single-level problems (26) and (28). Furthermore, all three problems have the same unique optimal solution $u^{*},y^{*}$ and $\sigma_{y}^{*}$.
The proof is provided in Appendix B.3. It is clear that (28) is in the form of (17). Thus, applying Theorem 1 leads to the following result.
Corollary 2.
Suppose $Q\succ 0,R\succ 0,\lambda_{y}>0$ and $\Gamma,\mathcal{U}$, and $\mathcal{Y}$ are compact and convex sets. If the optimal solution $(\sigma_{y}^{*},u^{*},y^{*})$ for (28) is an interior point of $\Gamma\times\mathcal{U}\times\mathcal{Y}$, then there exists $\lambda_{g}^{*}>0$, such that
$\min_{\begin{subarray}{c}g,\sigma_{y}\in\Gamma,\\ u\in\mathcal{U},y\in\mathcal{Y}\end{subarray}}$ $\|u\|_{R}^{2}+\|y\|_{Q}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2}$ (29)
$+\lambda_{g}\|Q_{\textnormal{c}}g\|_{2}$
$\mathrm{subject~to}$ $\hat{H}g=\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y},u,y),$
with $\hat{H}$ defined in (27), has the same optimal value and optimal solutions as (28) for all $\lambda_{g}>\lambda_{g}^{*}$.
We call (29) as Causality-based DDPC (C-DDPC). The new regularizer $\|Q_{\textnormal{c}}g\|_{2}$ penalizes both the violation of the row space constraint (i.e., $Q_{3}g$) and the usage of noncasual information (i.e., $Q^{*}g$).
Combining Proposition 2 and Corollary 2, we confirm that (29) is a relaxation of (24) under mild conditions.
Theorem 3.
Consider the bi-level DDPC (24) and the single-level C-DDPC (29). Suppose $Q\succ 0,R\succ 0,\lambda_{y}>0$ and $H_{\textnormal{d}}$ has full row rank with sufficiently large column number. Let $\Gamma,\mathcal{U}$ and $\mathcal{Y}$ are compact and convex sets and the optimal solution $(\sigma_{y}^{*},u^{*},y^{*})$ for (24) is an interior point of $\Gamma\times\mathcal{U}\times\mathcal{Y}$. For any positive $\lambda_{g}$, (29) is a convex relaxation of (24), that is:
  1. 1.
    (29) is a convex optimization problem;
  2. 2.
    any feasible $\sigma_{y},u,y$ for (24) is feasible to (29);
  3. 3.
    the optimal value of (29) is smaller than or equal to that of (24).
To our best knowledge, the single-level C-DDPC in (29) is new, and its relationship with the bi-level DDPC (24) has not been discussed before. The closest work on causality-based regularizer is [31], which regularizes on $\gamma$ in the $\gamma$-DDPC framework. Instead, our method (29) directly regularizes $g$, which can be more convenient to combine with $\|g\|_{1}$ for implicit trajectory selection.
4.3 Low-rank approximation and $l_{1}$ Norm
We finally consider the following bi-level problem
$\min_{\begin{subarray}{c}g,\sigma_{y}\in\Gamma,\\ u\in\mathcal{U},y\in\mathcal{Y}\end{subarray}}\quad$ $\|y\|_{Q}^{2}+\|u\|_{R}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2}$
$\mathrm{subject~to}\quad$ ${H}^{*}g=\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y},u,y),$ (30)
$\textrm{where}\;\;{H}^{*}\in\operatorname*{argmin}_{\tilde{H},K}\;\;\|\tilde{H}-H_{\mathrm{d}}\|_{F}^{2}$
[…]
This bi-level problem aims to approximate the data $H_{\textnormal{d}}$ with a low-rank matrix in (10a). It is clear that the unique optimal solution $H_{\textnormal{lr}}^{*}$ of the inner problem in (30) is
$H_{\textnormal{lr}}^{*}=\sum_{i=1}^{mL+n}\bar{\sigma}_{i}\bar{u}_{i}\bar{v}_{i}^{{\mathsf{T}}}$
where $\bar{\sigma}_{i},\bar{u}_{i}$ and $\bar{v}_{i}$ are obtained via the standard SVD decomposition $H_{\textnormal{d}}=\sum_{i=1}^{(m+p)L}\bar{\sigma}_{i}\bar{u}_{i}\bar{v}_{i}^{{\mathsf{T}}}$. Problem (30) can be equivalently formulated as
$\min_{g,\sigma_{y}\in\Gamma,u\in\mathcal{U},y\in\mathcal{Y}}$ $\|y\|_{Q}^{2}+\|u\|_{R}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2}$ (31)
$\mathrm{subject~to}$ $\tilde{H}_{\textnormal{lr}}^{*}g=\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y},u,y).$
We now fix $H=H_{\textnormal{d}}$ and add a $l_{1}$-norm constraint of $g$, leading to a single-level problem
$\min_{g,\sigma_{y}\in\Gamma,u\in\mathcal{U},y\in\mathcal{Y}}$ $\|y\|_{Q}^{2}+\|u\|_{R}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2}$ (32)
$\mathrm{subject~to}$ $H_{\textnormal{d}}g=\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y},u,y),$
$\|g\|_{1}\leq\alpha_{\textnormal{c}}.$
We have the following result.
Proposition 3.
Fix any data matrix $H_{\textnormal{d}}$, and suppose $\alpha_{\textnormal{c}}$ is sufficiently large. The optimal value of (32) is smaller than or equal to that of (31).
The proof is immediate from the fact that the optimal solution for (31) is feasible for (32) for sufficiently large $\alpha_{\textnormal{c}}$ as $\mathrm{range}(H^{*}_{\textnormal{lr}})\subseteq\mathrm{range}(H_{\textnormal{d}})$.
The constraint $\|g\|_{1}\leq\alpha_{\textnormal{c}}$ can be moved to the objective function without changing the optimal solution of (32) (but not the optimal value). We have the following result.
Proposition 4.
Suppose $Q\succ 0,R\succ 0,\lambda_{y}>0$, $\Gamma\times\mathcal{U}\times\mathcal{Y}$ is a convex set and the convex optimization problem (32) satisfies the Slater’s condition. Let $(\sigma_{y}^{*},u^{*},y^{*},g^{*})$ be an optimal solution for (32). There exists $\lambda_{g}^{*}\geq 0$, such that $(\sigma_{y}^{*},u^{*},y^{*},g^{*})$ is also an optimal solution for
$\min_{g,\sigma_{y}\in\Gamma,u\in\mathcal{U},y\in\mathcal{Y}}$ $\|y\|_{Q}^{2}+\|u\|_{R}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2}$ (33)
$+\lambda_{g}^{*}\|g\|_{1}$
$\mathrm{subject~to}$ $H_{\textnormal{d}}g=\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y},u,y).$
The proof is given in Appendix B.5. Propositions 3 and 4 imply that (33) is a relaxation of (30) with respect to the function $\|y\|_{Q}^{2}+\|u\|_{R}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2}$ under mild conditions.
Theorem 4.
Consider the bi-level problem (30) and the single-level problem (33). Suppose $Q\succ 0,R\succ 0$ and $\lambda_{y}>0$. Let $\Gamma,\mathcal{U}$ and $\mathcal{Y}$ be convex sets. There exists $\lambda_{g}^{*}\geq 0$, such that (33) is a convex relaxation of (30) with respect to the function $\|y\|_{Q}^{2}+\|u\|_{R}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2}$, that is:
  1. 1.
    (33) is a convex optimization problem;
  2. 2.
    any feasible $\sigma_{y},u,y$ for (30) is feasible to (33);
  3. 3.
    The value of $\|y\|_{Q}^{2}+\|u\|_{R}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2}$ evaluated at the optimal solution of (33) is smaller than or equal to that of (30).
A similar statement as Theorem 4 appears in [1, Theorem 4.8]. However, we emphasize that the relaxation is only true with respect to the function $\|y\|_{Q}+\|u\|_{R}+\lambda_{y}\|\sigma_{y}\|$, which is treated differently in [1]. The reason is that while the optimal solution of (33) is the same as (32), their objective functions are different, which generally leads to different optimal values.
5 Hybrid preprocessing and regularization of data for predictive control
The DDPC variants in Section 4 have modified the inner identification problem (10), instead of directly solving it. By simplifying the inner problem, we obtain single-level formulations that serve as convex relaxations for the modified bi-level DDPC (7); see Theorems 2, 3, and 4.
In this section, we propose an iterative algorithm that approximates the solution of the original inner problem (10) and leverages this approximate solution for the outer predictive control. In particular, given the offline input and output data (4), we consider that the input data $H_{u}:=\textnormal{col}(U_{\textnormal{P}},U_{\textnormal{F}})$ is accurate and contains no noise. Therefore, we focus on denoising the output trajectory, and problem (10) becomes
$\min_{\tilde{H}_{y},K}\quad$ $\|H_{y}-\tilde{H}_{y}\|_{F}$
$\mathrm{subject~to}\quad$ $\textrm{rank}(\tilde{H})=mL+n,$ (34a)
$\tilde{Y}_{\textnormal{F}}=K\ \textnormal{col}(U_{\textnormal{P}},\tilde{Y}_{\textnormal{P}},U_{\textnormal{F}}),$ (34b)
$K=\begin{bmatrix}K_{p}&K_{f}\end{bmatrix},\ K_{f}\in\mathcal{L},$ (34c)
$\tilde{H}_{y}\in\mathcal{H},$ (34d)
where $\tilde{H}_{y},\tilde{H}$ denote $\textnormal{col}(\tilde{Y}_{\textnormal{P}},\tilde{Y}_{\textnormal{F}})$,and $\textnormal{col}(U_{\textnormal{P}},\tilde{Y}_{\textnormal{P}},U_{\textnormal{F}},\tilde{Y}_{\textnormal{F}})$, respectively. Compared with (10), the decision variables in (34) become $\tilde{H}_{y}$ and $K$, as we only denoise the output trajectory. Still, problem (34) is difficult to solve due to the interplay between (34a) to (34c). Without (34b), (34c), it is a structured low-rank approximation (SLRA) problem [46, 47]. We here adapt an iterative SLRA algorithm in [47] to get the approximation solution to (34).
5.1 Sequential optimization
Our key idea is to adopt a sequential optimization strategy that addresses the constraints in (34) one at a time. The detailed procedure is described below.
Step 1: Low-rank approximation. Since the input data $u_{\textnormal{d}}$ has no noise and satisfies the persistent excitation, we have $\textrm{rank}(H_{u})=mL$. In contrast, the output data $y_{\textnormal{d}}$ contains “variance” noise and “bias” error, and thus the rank of the raw data $H$ (i.e., $\textnormal{col}(H_{u},H_{y})$) is larger than $mL+n$. We first consider the constraint (34a) that approximates $H$ with a low-rank matrix, which is
$\min_{\tilde{H}_{y}}$ $\|H_{y}-\tilde{H}_{y}\|_{F}$ (35)
$\mathrm{subject~to}$ $\textrm{rank}(\tilde{H})=mL+n.$
This problem is slightly different from the standard low-rank approximation, as we do not change the input data $U_{\textnormal{P}}$ and $U_{\textnormal{F}}$. Still, we can get an analytical solution of (35) via singular value decomposition.
Proposition 5.
Consider problem (35), where the input data $u_{\textnormal{d}}$ has no noise and satisfies the persistent excitation. Let $\Pi_{2}=H_{u}^{\dagger}H_{u}$ be the orthogonal projector onto the row space of $H_{u}$, and denote SVD for the component of $H_{y}$ in the null space of $H_{u}$ as $H_{y}(I-\Pi_{2})=\sum_{i=1}^{pL}\bar{\sigma}_{i}\bar{u}_{i}\bar{v}_{i}^{{\mathsf{T}}}$. Then, the optimal solution $H_{y_{1}}$ to (35) is given by
$H_{y_{1}}:=H_{y}\Pi_{2}+\sum_{i=1}^{n}\bar{\sigma}_{i}\bar{u}_{i}\bar{v}_{i}^{{\mathsf{T}}}.$ (36)
Since we have $\mathrm{rank}(H_{u})=mL$, the key insight for Proposition 5 is to ensure the part of $H_{y}$ in the null space of $H_{u}$ has rank $n$. Specifically, we first divide $H_{y}$ into two parts that are $H_{y}\Pi_{2}$ in the row space of $H_{u}$ and $H_{y}(I-\Pi_{2})$ in the null space of $H_{u}$. We then perform an SVD of $H_{y}(I-\Pi_{2})$ to estimate its rank-$n$ approximation, and finally combine it with the component of $H_{y}$ in the row space of $H_{u}$ as in (36).
For notational simplicity, we denote the mapping from the data $H_{y}$ to the optimal solution $H_{y_{1}}$ of (35) as $\Pi_{\textnormal{L}}$.
Step 2: Hankel structure: We then project $H_{y_{1}}$ to a Hankel matrix set via averaging skew-diagonal elements [47] and represent the projector and the resulting Hankel matrix as $\Pi_{\textnormal{H}}$ and $H_{y_{2}}$.
Step 3: Causality guarantee: We finally use the Hankel approximation $H_{y_{2}}$ to form the problem
$\min_{\tilde{H}_{y},K}$ $\|H_{y_{2}}-\tilde{H}_{y}\|_{F}$ (37)
$\mathrm{subject~to}$ $\tilde{Y}_{\textnormal{F}}=K\ \textnormal{col}(U_{\textnormal{P}},Y_{\textnormal{P}_{2}},U_{\textnormal{F}}),$
$K=\begin{bmatrix}K_{p}&K_{f}\end{bmatrix},\ K_{f}\in\mathcal{L},$
which tackles constraints (34b), (34c) and $\textnormal{col}(Y_{\textnormal{P}_{2}},Y_{\textnormal{F}_{2}})\!:=\!H_{y_{2}}$. Problem (37) also has an analytical solution as shown in (25), and we denote the mapping from $H_{y_{2}}$ to the optimal solution $H_{y_{3}}$ as $\Pi_{\textnormal{C}}$. We note that the analytical solution of (37) is derived based on the fact that $\textnormal{col}(U_{\textnormal{P}},Y_{\textnormal{P}_{2}},U_{\textnormal{F}},Y_{\textnormal{F}_{2}})$ has full row rank, which is generally true after the Hankel-structure approximation.
We repeat Steps 1 - 3 iteratively, and this process is listed in Algorithm 1. The resulting matrix $H_{y}^{*}$ from Algorithm 1 is partitioned as $\textnormal{col}(Y_{\textnormal{P}}^{*},Y_{\textnormal{F}}^{*})$, and we form a new Hankel matrix $H_{\textnormal{op}}^{*}=\textnormal{col}(U_{\textnormal{P}},Y_{\textnormal{P}}^{*},U_{\textnormal{F}},Y_{\textnormal{F}}^{*})$. We call $H_{\textnormal{op}}^{*}$ an approximated optimal linear representation, as it directly tackles the inner identification problem without relaxation.
Algorithm 1 Iterative Construction for Approximatedly Optimal Trajectory Library
[…]
3:repeat
4:  $H_{y_{1}}\leftarrow\Pi_{\textnormal{L}}(H_{y_{3}})$  (Low-rank approx);
5:  $H_{y_{2}}\!\leftarrow\!\!\Pi_{\textnormal{H}}(H_{y_{1}})$  (Hankel proj);
6:  $H_{y_{3}}\!\leftarrow\!\!\Pi_{\textnormal{C}}(H_{y_{2}})$  (Causality proj);
7:until $\|H_{y_{2}}-H_{y_{3}}\|_{F}\leq\epsilon\|H_{y_{3}}\|_{F}$
[…]
5.2 Data-driven predictive control with approximated linear data-driven representation
We use the approximated optimal linear model $H_{\textnormal{op}}^{*}$ from Algorithm 1 as the predictor in data-driven predictive control, which leads to
$\min_{\begin{subarray}{c}g,\sigma_{y}\in\Gamma,\\ u\in\mathcal{U},y\in\mathcal{Y}\end{subarray}}$ $\|u\|_{R}^{2}+\|y\|_{Q}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2}+\lambda_{1}\|g\|_{1}$ (38)
$\mathrm{subject~to}$ $H_{\textnormal{op}}^{*}g\!=\!\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y},u,y),$
where the $l_{1}$-norm regularizer is to implicitly select the trajectory in $H_{\textnormal{op}}^{*}$. We call this formulation (38) as Optimality-based DDPC (O-DDPC).
Note that utilizing SVD and projecting the resulting matrix to the set of Hankel matrices iteratively has been well-studied; see, e.g., [46, 47]. The convergence for the iterative algorithm using Hankel projection only is discussed in [48]. However, in Algorithm 1, we require the extra separation step to divide $H_{y}$ into the row and null spaces of $H_{u}$ and the causality projection. While our numerical simulations indicate that Algorithm 1 converges in practice, establishing a formal theoretical proof remains an open problem, which we leave for future work.
Remark 3 (LTI system characterizations).
The constraints (34a)-(34c) are necessary for the Hankel matrix $H$ to come from an LTI system, but they are not sufficient. As discussed in [1], the parameter $K$ needs to satisfy additional structure requirements to ensure the existence of the corresponding matrix parameters $A,B,C,D$ in (1). When data is collected from nonlinear systems, making the data more structured may introduce a larger bias error. Our proposed O-DDPC in (38) preserves the linear, low-rank structure, and causality without requiring the predictor $H_{\textnormal{op}}^{*}$ to exactly correspond to an LTI system. This hybrid preprocessing and regularization imposes more system structure but also gives freedom may allow (38) to select model complexity implicitly, thus improving empirical control performance.
6 Numerical experiments
We compare the numerical performance [^1] Our code is available at https://github.com/soc-ucsd/Convex-Approximation-for-DeePC. for the DDPC variants in Sections 4 and 5, including: 1) Linearity-based DDPC (23) (L-DDPC), 2) Causality-based DDPC (29) (C-DDPC) and 3) Optimality-based DDPC (38) (O-DDPC). We also include results from the Hankel-form SPC (SPC) in (21) and the indirect DDPC using system identification for comparison. In particular, we used the N4SID [49] for the system ID. We do not consider (33) individually, as an $l_{1}$-norm regularizer is added for all DDPC variants and the Hankel-form SPC.
We perform two sets of experiments: 1) an LTI system with noisy measurement [17] and 2) another nonlinear system with noise-free measurement [1] (corresponding to “variance” error and “bias” error, per [1]). Our results show the superior performance of O-DDPC in both systems. We also see that the indirect method performs better under the “variance” error, while the direct DDPC variants have superior performance under the “bias” error. This finding is consistent with [1].
6.1 An LTI system
Experiment setup. We first consider an LTI system from [17], which is a triple-mass-spring system with $n=8$ states, $m=2$ inputs (two stepper motors), and $p=3$ outputs (disc angles). We consider data collection with additive Gaussian measurement noises $\omega\sim\mathcal{N}(0,\sigma I)$. We utilize the data sets with various sizes and noise levels that are $T=400,600,800$ and $\sigma$ from $0.02$ to $0.1$ in increments of $0.02$. The prediction horizon and the initial sequence are chosen as $N=40$ and $T_{\textnormal{ini}}=4$, respectively. We choose $Q=I$, $R=0.1I,\mathcal{U}=[-0.7,0.7]$ and $\Gamma,\mathcal{Y}=\mathbb{R}^{3}$ unless otherwise specified.
Numerical results. We here compare the realized control cost for different DDPC variants with various sizes of data sets and noise levels. Since the performance of DDPC variants depends on the pre-collected data set, all realized control costs for different DDPC variants are averaged over 100 pre-collected data sets.
(a) T = 400
(b) T = 600
(c) T = 800
Figure 4: Comparison of realized control cost of different variants of DDPC with varying measurement noises. (a), (b) and (c) illustrate the control performance of data-driven controllers with different pre-collected trajectory lengths. Across all experiments, the proposed O-DDPC (38) is comparable to the system ID approach, which provides the best control performance among all DDPC methods.
The numerical results are shown in Fig. 4. As expected, the control performance of all data-driven controllers deteriorates with increasing noise levels. Among the DDPC-based methods, the proposed O-DDPC (38) achieves the best realized control costs and exhibits the strongest noise robustness. This improvement may come from its iterative low-rank approximation, causality projection, and Hankel structure projection (see Algorithm 1), which together act as an effective noise filter. This reduces the influence of variance noises in LTI systems.
Table 1: Realized Control Cost and with different pre-collected trajectory lengths with $\sigma=0.1$; The ground truth value is $277.2487$, and the increase ratio is shown in parentheses.
$T=400$ $T=600$ $T=800$
L-DDPC 377.51 ($36.16\%$) 361.36 ($30.34\%$) 349.10 ($25.92\%$)
SPC 368.42 ($32.88\%$) 356.60 ($28.62\%$) 346.56 ($25.00\%$)
C-DDPC 364.51 ($31.48\%$) 351.43 ($26.76\%$) 340.39 ($22.77\%$)
O-DDPC 300.77 ($\mathbf{8.48\%}$) 296.59 ($\mathbf{6.98\%}$) 296.31 ($\mathbf{6.88\%}$)
SysID 289.25 ($4.33\%$) 284.64 ($2.67\%$) 283.02 ($2.08\%$)
Table 1 lists the realized control costs at noise level $\sigma=0.1$ for different pre-collected trajectory lengths and control approaches. We also report their percentage increases relative to the ground-truth cost computed with noise-free data from (6). The results show a clear performance ordering: $\texttt{L-DDPC}>\texttt{SPC}>\texttt{C-DDPC}>\texttt{O-DDPC}>\textrm{System ID}$. For the LTI system with noisy data, the inner problem in (7) enforces LTI structure in the data-driven representation, which in turn improves noise rejection in the outer predictive control problem (7). Consequently, methods that impose more constraints on the inner problem tend to yield more structured representations and better performance. Notably, the increase in realized control cost for O-DDPC is approximately $7\%$, significantly outperforming the other DDPC variants.
6.2 Nonlinear system
We here consider a nonlinear Lotka-Volterra system [1]
$\dot{x}=\begin{bmatrix}\dot{x}_{1}\\ \dot{x}_{2}\end{bmatrix}=\begin{bmatrix}ax_{1}-bx_{1}x_{2}\\ dx_{1}x_{2}-cx_{2}+u\end{bmatrix}$, (39)
where $x_{1},x_{2}$ denote prey and predator populations and $u$ is the control input. We used $a=c=0.5,b=0.025,d=0.005$ in our experiments. We first linearize the nonlinear system (39) around the equilibrium $(\bar{u},\bar{x}_{1},\bar{x_{2}})=(0,\frac{c}{d},\frac{a}{b})$. Then, after discretization, we obtain a linear system as
$\hat{x}(k+1)$ $=f_{\text{linear}}(\hat{x}(k),\hat{u}(k))$
$=\begin{bmatrix}\hat{x}_{1}(k)+\Delta t(-b\bar{x}_{1}\hat{x}_{2}(k))\\ \hat{x}_{2}(k)+\Delta t(d\bar{x}_{2}\hat{x}_{1}(k)+\hat{u}(k))\end{bmatrix}$
where $\Delta t=0.1$ is the time step for discretization. We discretize the nonlinear system in the error state space
$\hat{x}(k+1)$ $=f_{\text{nonlinear}}(\hat{x}(k),\hat{u}(k))$
$=\begin{bmatrix}\hat{x}_{1}(k)+\Delta t(a(\hat{x}_{1}(k)+\bar{x_{1}})\\ -b(\hat{x}_{1}(k)+\bar{x_{1}})(\hat{x}_{2}(k)+\bar{x_{2}}))\\ \hat{x}_{2}(k)+\Delta t(d(\hat{x}_{1}(k)+\bar{x_{1}})(\hat{x}_{2}(k)+\bar{x_{2}})\\ -c(\hat{x}_{2}(k)+\bar{x_{2}})+\hat{u}(k))\end{bmatrix}$.
We then construct systems with various nonlinearity by interpolating between $f_{\textrm{linear}}$ and $f_{\textrm{nonlinear}}$ that is
$\hat{x}(k+1)=$ $\epsilon\cdot f_{\text{linear}}(\hat{x}(k),\hat{u}(k))$
$\quad+(1-\epsilon)\cdot f_{\text{nonlinear}}(\hat{x}(k),\hat{u}(k)).$
The length of the pre-collected trajectory is $T=400$. The prediction horizon and initial sequence are set as $N=60$ and $T_{\textnormal{ini}}=4$, respectively. We choose $Q=I,R=0.5I$ and $\mathcal{U}=[-20,20]$. The parameters for each controller are the same as those in Section 6.1.
Comparison of direct/indirect methods. We compare the realized control costs for the indirect system ID approach and different DDPC variants on systems with varying degrees of nonlinearity. Model orders are chosen to be 8 and 4 for O-DDPC in Algorithm 1 and N4SID, respectively. Similar to Section 6.1, we average the realized control costs over 100 pre-collected trajectories. We note that the identified model from N4SID is often ill-conditioned when the nonlinearity is high, which caused numerical issues in solving (2) in some of our experiments. We discard these ill-conditioned solutions when computing the average performance for the indirect system ID approach.
Both direct and indirect approaches perform well when the nonlinearity is low (i.e., $\epsilon\in[0.8,1])$, as shown in Fig. 5. However, the cost for the indirect method significantly increases with higher nonlinearity, while the performance of direct methods remains relatively consistent. The superior performance of direct data-driven methods is consistent with experimental observations from [1]. The indirect system ID method projects the noisy data on a fixed linear model, which induces “bias” error due to selecting a wrong model class; on the other hand, the complexity of the LTI system is regularized but not specified in direct methods, which provides more flexibility and leads to superior performance for controlling nonlinear systems in our experiments.
Figure 5: Comparison of realized cost for system ID and direct DDPC variants for the nonlinear system in Section 6.2. The System ID approach has the worst performance due to a wrong model class.
Comparison of DDPC variants. We then compare the performance of different DDPC variants. The results are shown in Fig. 6. The C-DDPC and O-DDPC with more structured data-driven representation outperform L-DDPC and SPC which only relax or tackle the linearity requirement without considering the causality and the Hankel structure. Furthermore, among all direct data-driven approaches, O-DDPC performs the best for both the nonlinear system (bias error) and the LTI system with measurement noise (variance noise). These numerical results suggest that we might obtain additional benefits when employing appropriate techniques from system ID to pre-process the trajectory library of nonlinear systems.
Figure 6: Comparison of realized control cost of different DDPC variants for systems with varying nonlinearity. Across all experiments, the proposed O-DDPC (38) provides the best control performance among all DDPC methods.
7 Conclusion
In this paper, we have analyzed the role of regularizations in direct and indirect data-driven control via a bi-level optimization framework. We prove that, after dropping some inner constraints, the bi-level optimization problem can be transformed into a single-level convex optimization problem. Moreover, regularizers developed from those single-level problems are exact penalty functions under certain conditions. Thus, using regularizers is a further convex relaxation with respect to the original bi-level problem (after dropping constraints). Moreover, we have proposed a new variant, called O-DDPC, which uses an iterative algorithm to obtain a novel data-driven control approach. Numerical simulations have demonstrated the superior performance of O-DDPC (38) with the more structured predictor. Future directions include investigating the closed-loop performance across various DDPC variants, and validating the proposed O-DDPC approach through real-world experiments.
Appendix A Exact penalty for quadratic optimization
The proof for Theorem 1 is divided into two parts:
  1. 1.
    We transform (17) and (18) to two new optimization problems, where the optimal solutions $x_{1}^{*}$ remains the same.
  2. 2.
    We next establish an exact penalty such that new formulations have the same optimal solutions with a finite regularization parameter.
We first change the variables in (17) and (18) to facilitate the penalty analysis. The key idea is to construct a new decision variable $\bar{x}_{2}$ such that $\bar{x}_{2}=Dx_{2}$ with a free decision variable $x_{3}$. We consider two problems
$\min_{x_{1}\in\mathcal{X},\bar{x}_{2}\in\mathcal{D},x_{3}}$ $x_{1}^{{\mathsf{T}}}Mx_{1}$ (40)
$\mathrm{subject~to}$ $A_{1}x_{1}+A_{2}D^{\dagger}\bar{x}_{2}+A_{2}D^{\bot}x_{3}=b,$
$\|\bar{x}_{2}\|_{p}=0,$
and its penalized version
$\min_{x_{1}\in\mathcal{X},\bar{x}_{2}\in\mathcal{D},x_{3}}$ $x_{1}^{{\mathsf{T}}}Mx_{1}+\lambda_{w}\|\bar{x}_{2}\|_{p}$ (41)
$\mathrm{subject~to}$ $A_{1}x_{1}+A_{2}D^{\dagger}\bar{x}_{2}+A_{2}D^{\bot}x_{3}=b,$
where columns of $D^{\bot}$ form a basis for the null space of $D$ and $\mathcal{D}:=\mathrm{Im}(D)$. We have the following equivalence among optimization problems (17), (18), (40) and (41)
Proposition 6.
Problem (17) and (40) share the same optimal solution $x_{1}^{*}$, and problem (18) and (41) also share the same optimal solution $x_{1}^{*}$ for any $\lambda_{w}\geq 0$.
Proof.
We first show the equivalence of (17) and (40):
  1. 1.
    (17) and (40) have the same feasible region: if $x_{1},x_{2}$ is feasible to (17), we can construct $\bar{x}_{2}$ and $x_{3}$ such that $x_{1},\bar{x}_{2},x_{3}$ is feasible to (40). On the other hand, given any feasible solution $x_{1},\bar{x}_{2},x_{3}$, there exists $x_{2}$ such that $x_{1},x_{2}$ is feasible to (17).
  2. 2.
    The (17) and (40) have the same cost function over their feasible region.
Let $x_{1},x_{2}$ be a feasible solution to (17). We construct $\bar{x}_{2}$ as $\bar{x}_{2}=Dx_{2}$ and $x_{3}$ such that $D^{\bot}x_{3}=x_{2}-D^{\dagger}\bar{x}_{2}=x_{2}-D^{\dagger}Dx_{2}.$ The existence of $x_{3}$ is guaranteed as $D^{\dagger}Dx_{2}$ represents the projection of $x_{2}$ onto the row space of $D$, so that the residual must be included in the range space of $D^{\bot}$. Thus, we have $\|\bar{x}_{2}\|_{p}=\|Dx_{2}\|_{p}=0$ and
$A_{1}x_{1}+A_{2}D^{\dagger}\bar{x}_{2}+A_{2}D^{\bot}x_{3}=A_{1}x_{1}+A_{2}x_{2}=b,$
which means $x_{1},\bar{x}_{2},x_{3}$ is a feasible solution of (40).
We then let $x_{1},\bar{x}_{2},x_{3}$ is be feasible solution of (40). We now let $x_{2}=D^{\dagger}\bar{x}_{2}+D^{\bot}x_{3}$, and write
$\|Dx_{2}\|_{p}=\|DD^{\dagger}\bar{x}_{2}+DD^{\bot}x_{3}\|_{p}=\|\bar{x}_{2}\|_{p}=0,$
$A_{1}x_{1}+A_{2}x_{2}=A_{1}x_{1}+A_{2}D^{\dagger}\bar{x}_{2}+A_{2}D^{\bot}x_{3}=b,$
This shows $x_{1},x_{2}$ is a feasible solution for (17). Similarly, we can prove the equivalence of (18) and (41).  ∎
We next prove that $\|\bar{x}_{2}\|_{p}$ is an exactly penalty function for (40). Consider a perturbed version of (40):
$\min_{x_{1}\in\mathcal{X},\bar{x}_{2}\in\mathcal{D},x_{3}}$ $x_{1}^{{\mathsf{T}}}Mx_{1}$ (42)
$\mathrm{subject~to}$ $A_{1}x_{1}+A_{2}T^{\dagger}\bar{x}_{2}+A_{2}T^{\bot}x_{3}=b,$
$\|\bar{x}_{2}\|_{p}=\epsilon.$
We will use (42) to establish the following proposition.
Proposition 7.
Suppose $\mathcal{X}$ is a convex set, $M$ is positive semidefinite and $x_{1}^{{\mathsf{T}}}Mx_{1}$ is locally Lipchitz continuous with Lipchitz constant $L$ with respect to $\|\cdot\|_{p}$ over $\mathcal{X}$. If there exists an optimal solution $(x_{1}^{*},\bar{x}_{2}^{*},x_{3}^{*})$ for (40) such that $x_{1}^{*}$ is an interior point of $\mathcal{X}$, there exists $\lambda_{w}^{*}>0$, such that (40) and (41) have same optimal solutions for all $\lambda_{w}>\lambda_{w}^{*}$.
Proof.
The key idea is to prove that the optimal solution $(x_{1}^{*},\bar{x}_{2}^{*},x_{3}^{*})$ of (40) is a local minimum of (41) via the usage of (42). This optimal solution also becomes the global minimum since (41) is a convex problem. This is inspired by the notion of partial calmness [44].
  1. 1.
    Problem (42) permits a certain degree of “violation” of the constraint $\|\bar{x}_{2}\|_{p}=0$. We first show that, in a local neighborhood of $(x_{1}^{*},\bar{x}_{2}^{*},x_{3}^{*})$, such violations provide no benefit once incorporated into the objective function with an appropriate weighting parameter. This implies that $(x_{1}^{*},\bar{x}_{2}^{*},x_{3}^{*})$ is a local optimal solution of (42).
  2. 2.
    Next, we establish that there exists a local region around $(x_{1}^{*},\bar{x}_{2}^{*},x_{3}^{*})$ in which every feasible solution of (41) is also feasible for (42). Combining this fact with the result of Step 1, we conclude that $(x_{1}^{*},\bar{x}_{2}^{*},x_{3}^{*})$ is a local optimal solution of (41). It is thus globally optimal due to the convexity of (41)
  3. 3.
    Finally, we demonstrate that all optimal solutions of (41) are also optimal solutions of (40).
Let $A_{3}:=[A_{1},A_{2}D^{\bot}]$ and $P_{2}:=\textnormal{col}(I,\mathbb{0})$. As $x_{1}^{*}\in\mathcal{X}^{\circ}$, there exists $\delta_{r}>0$ such that $\mathcal{B}_{\delta_{r}}(x_{1}^{*})\subseteq\mathcal{X}$. Let $\frac{\delta_{r}}{\sigma_{r}+\sigma_{p}}>\delta>0$ where $\sigma_{r}$ is the induced $p$-norm of $P_{2}A_{3}^{\dagger}A_{2}D^{\dagger}$ and $\sigma_{p}$ is the constant such that $\|v\|_{p}\leq\sigma_{p}\|v\|_{2}$ for any vector $v$ with the same dimension of $x_{1}$. For any $\epsilon\in[0,\delta)$, let $(x_{1},\bar{x}_{2},x_{3})\in\mathcal{B}_{\delta}(x_{1}^{*},\bar{x}_{2}^{*},x_{3}^{*})$ be a feasible solution for (42). We first show that there exists $\mu$, such that
$x_{1}^{{\mathsf{T}}}Mx_{1}+\mu\epsilon\geq x_{1}^{*{{\mathsf{T}}}}Mx_{1}^{*}.$ (43)
We can represent $\textnormal{col}(x_{1},x_{3})$ as
$\textnormal{col}(x_{1},x_{3})=A_{3}^{\dagger}(b-A_{2}D^{\dagger}\bar{x}_{2})+v_{1},$
where $v_{1}\in\mathrm{Null}(A_{3})$. A feasible solution $(\tilde{x}_{1},\bar{x}_{2}^{*},\tilde{x}_{3})$ for (40) can be constructed as
$\textnormal{col}(\tilde{x}_{1},\tilde{x}_{3})=A_{3}^{\dagger}(b-A_{2}D^{\dagger}\bar{x}_{2}^{*})+v_{1}.$
We then demonstrate $\tilde{x}_{1}\in\mathcal{X}$ via deriving the difference between $x_{1}$ and $\tilde{x}_{1}$. We have
$\|x_{1}-\tilde{x}_{1}\|_{p}=\|P_{2}A_{3}^{\dagger}A_{2}D^{\dagger}\bar{x}_{2}\|_{p}\leq\sigma_{r}\epsilon\leq\sigma_{r}\delta.$
We then obtain
$\|x_{1}^{*}-\tilde{x}_{1}\|_{p}\leq\|x_{1}^{*}-x_{1}\|_{p}+\|x_{1}-\tilde{x}_{1}\|_{p}\leq(\sigma_{r}+\sigma_{p})\delta\leq\delta_{r},$
which implies $\tilde{x}_{1}\in\mathcal{B}_{\delta_{r}}(x^{*})\subseteq\mathcal{X}$. From the optimality condition of (40), we have
$x_{1}^{{\mathsf{T}}}Mx_{1}-x^{*{{\mathsf{T}}}}Mx^{*}\geq$ $x_{1}^{{\mathsf{T}}}Mx_{1}-\tilde{x}_{1}^{{\mathsf{T}}}M\tilde{x}_{1}$
$\geq$ $-L\|x_{1}-\tilde{x}_{1}\|_{p}\geq-L\sigma_{r}\epsilon,$
which means $x_{1}^{{\mathsf{T}}}Mx_{1}+L\sigma_{r}\epsilon\geq x_{1}^{*{{\mathsf{T}}}}Mx_{1}^{*}.$ Thus, we can let $\mu\geq L\sigma_{r}$ so that the inequality (43) is satisfied.
We then prove that the optimal solution $(x_{1}^{*},\bar{x}_{2}^{*},x_{3}^{*})$ of (40) is a local minima of (41). We can let $\lambda_{w}>\lambda_{w}^{*}\geq L\sigma$. Since $f(x):=\|x\|_{p}$ is a continuous function, there exists $\delta_{x}>0$ such that $\|\bar{x}_{2}\|_{p}<\delta$ for any $\bar{x}_{2}\in\mathcal{B}_{\delta_{x}}(\bar{x}_{2}^{*})$. Let $\delta^{*}:=\min(\delta_{x},\delta)$. For any feasible solution $(x_{1},\bar{x}_{2},x_{3})\in\mathcal{B}_{\delta^{*}}(x_{1}^{*},\bar{x}_{2}^{*},x_{3}^{*})$ of (41), there exists $\epsilon\in[0,\delta)$ such that it is feasible to (42). Thus, from the inequality (43), we have for any $(x_{1},\bar{x}_{2},x_{3})\in\mathcal{B}_{\delta^{*}}(x_{1}^{*},\bar{x}_{2}^{*},x_{3}^{*})$,
$x_{1}^{{\mathsf{T}}}Mx_{1}+\lambda_{w}\|\bar{x}_{2}\|_{p}\geq x_{1}^{*{{\mathsf{T}}}}Mx_{1}^{*},$
which means $(x_{1}^{*},\bar{x}_{2}^{*},x_{3}^{*})$ is a local minima (as well as the global minima) of (41).
It is now clear that all optimal solutions of (40) are optimal to (41). Next, suppose $(x_{1w}^{*},\bar{x}_{2w}^{*},x_{3w}^{*})$ is an optimal solution of (41) and recall that $\lambda_{w}>\lambda_{w}^{*}\geq L\sigma_{r}$. We have
$x_{1w}^{*{{\mathsf{T}}}}Mx_{1w}^{*}+\lambda_{w}\|\bar{x}_{2w}^{*}\|_{p}=x_{1}^{*{{\mathsf{T}}}}Mx_{1}^{*}\leq x_{1w}^{*{{\mathsf{T}}}}Mx_{1w}^{*}+\lambda_{w}^{*}\|\bar{x}_{2w}^{*}\|_{p},$
which leads to $(\mu-\bar{\mu})\|\bar{x}_{2w}^{*}\|_{2}\leq 0$ and implies $\|\bar{x}_{2w}^{*}\|_{2}=0$. Thus, $(x_{1w}^{*},\bar{x}_{2w}^{*},x_{3w}^{*})$ is also feasible for (40) and is its optimal solution.  ∎
We combine Propositions 6 and 7 to establish Theorem 1.
Proof of Theorem 1: From Proposition 6, the optimal solution $x_{1}^{*}$ to (17) and (40) is the same. Furthermore, the assumption that $(x_{1}^{*},x_{2}^{*})$ is an optimal solution with $x_{1}^{*}\in\mathcal{X}^{\circ}$ for (17) implies there exists an optimal solution $(x_{1}^{*},\bar{x}_{2}^{*},x_{3}^{*})$ and $x_{1}^{*}\in\mathcal{X}^{\circ}$ for (40). Thus, using Proposition 7, there exists $\lambda_{w}^{*}$ such that the optimal solution sets $x_{1}^{*}$ for (40) and (41) are equivalent for all $\lambda_{w}>\lambda_{w}^{*}$. That means those of (17) and (41) are the same with $\lambda_{w}^{*}$. Since we also have the optimal solutions of $x_{1}^{*}$ are the same for (18) and (41), this illustrates (17) and (18) have the same optimal solution set for $x_{1}^{*}$ with $\lambda_{w}^{*}$.
From Propositions 6 and 7, the optimal values of (17) and (18) are the same. As any optimal solutions $(x_{1}^{*},x_{2}^{*})$ for (17) are feasible for (18), substituting $(x_{1}^{*},x_{2}^{*})$ into the objective function of (18) we have
$x_{1}^{*{{\mathsf{T}}}}Mx_{1}^{*}+\lambda_{w}\|Dx_{2}^{*}\|_{p}=x_{1}^{*{{\mathsf{T}}}}Mx_{1}^{*},$
which means $(x_{1}^{*},x_{2}^{*})$ is an optimal solution for (18). Then, suppose $(x_{1}^{*},x_{2}^{*})$ is an optimal solution for (18). Since $x_{1}^{*}$ is also an optimal solution for (17) and their optimal values are the same, we have
$x_{1}^{*{{\mathsf{T}}}}Mx_{1}^{*}+\lambda_{w}\|Dx_{2}^{*}\|_{p}=x_{1}^{*{{\mathsf{T}}}}Mx_{1}^{*}\ \Rightarrow\ \|Dx_{2}^{*}\|_{p}=0.$
Thus, $(x_{1}^{*},x_{2}^{*})$ is an optimal solution for (17). This completes the proof.
Appendix B Technical proofs
B.1 Relation with the classcial SPC
The classical SPC is of the following form
$\min_{\sigma_{y}\in\Gamma,u\in\mathcal{U},y\in\mathcal{Y}}$ $\|y\|_{Q}^{2}+\|u\|_{R}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2}$ (44)
$\mathrm{subject~to}$ $y=Y_{\textnormal{F}}\begin{bmatrix}U_{\textnormal{P}}\\ Y_{p}\\ U_{\textnormal{F}}\end{bmatrix}^{\dagger}\begin{bmatrix}u_{\textnormal{ini}}\\ y_{\textnormal{ini}}+\sigma_{y}\\ u\\ \end{bmatrix}$,
which does not have the variable $g$. We can establish the following equivalence.
Proposition 8.
If $Q\!\succ\!0,R\!\succ\!0$ and $H_{1}\!=\!\textnormal{col}(U_{\textnormal{P}},Y_{\textnormal{P}},U_{\textnormal{F}})$ in (20) has full row rank, then (21) and (44) have the same optimal solution $u^{*},y^{*}$ and $\sigma_{y}^{*}$, $\forall\lambda_{y}>0$.
Our proof is divided into two main parts:
  1. 1.
    When $H_{1}=\textnormal{col}(U_{\textnormal{P}},Y_{\textnormal{P}},U_{\textnormal{F}})$ has full row rank, we show that (21) and (44) have the same feasible region: if $\sigma_{y},u,y,g$ is feasible to (21), then the same $\sigma_{y},u,y$ is also feasible for (44). Conversely, given any feasible solution $\sigma_{y},u,y$ to (44), there exists a vector $g$ such that $\sigma_{y},u,y,g$ is feasible to (21).
  2. 2.
    The (21) and (44) have the same cost function in terms of $u,y,\sigma_{y}$.
Combining the two properties above with the fact that the cost function in (44) is strongly convex, we conclude (21) and (44) have the same unique optimal solution for decision variables $\sigma_{y},u$ and $y$.
The property 2 above is obvious. We prove the property 1 below. Let us first decompose
$Y_{\textnormal{F}}=M+M^{\bot},$ (45)
where $M$ (i.e., $Y_{\textnormal{F}}\Pi_{1}$) is the (row space) orthogonal projection of $Y_{\textnormal{F}}$ on the row space of $H_{1}$ and $M^{\bot}$ (i.e., $Y_{\textnormal{F}}(I-\Pi_{1})$) is the rest part of $Y_{\textnormal{F}}$ in the null space of $H_{1}$. Since $H_{1}$ has full row rank, thanks to the property of Moore–Penrose inverse, we have $H_{1}H_{1}^{\dagger}=I$ and the range space of $H_{1}^{\dagger}$ is the same as the row space of $H_{1}$, which means $M^{\bot}H_{1}^{\dagger}=0$.
We assume $u_{1},y_{1},\sigma_{y_{1}},g_{1}$ is a feasible solution for (21). Then, without loss of generality, the vector $g_{1}$ can be represented as
$g_{1}=H_{1}^{\dagger}\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y_{1}},u_{1})+\hat{g}$
where $\hat{g}$ is a vector in the null space of $H_{1}$. We have $M\hat{g}=0$ since $H_{1}\hat{g}=0$, and $Y_{\textnormal{F}}H_{1}^{\dagger}=(M+M^{\bot})H_{1}^{\dagger}=MH_{1}^{\dagger}$ because $M^{\bot}H_{1}^{\dagger}=0$. Thus, from the equality constrain in (21), the vector $y_{1}$ satisfies
$y_{1}$ $=Mg_{1}=MH_{1}^{\dagger}\begin{bmatrix}u_{\textnormal{ini}}\\ y_{\textnormal{ini}}+\sigma_{y_{1}}\\ u_{1}\end{bmatrix}+M\hat{g}$ (46)
$=MH_{1}^{\dagger}\begin{bmatrix}u_{\textnormal{ini}}\\ y_{\textnormal{ini}}+\sigma_{y_{1}}\\ u_{1}\end{bmatrix}=Y_{\textnormal{F}}H_{1}^{\dagger}\begin{bmatrix}u_{\textnormal{ini}}\\ y_{\textnormal{ini}}+\sigma_{y_{1}}\\ u_{1}\end{bmatrix}$,
[…]
We next assume $u_{1},y_{1},\sigma_{y_{1}}$ is a feasible solution for (44). Substituting the orthonormal decomposition (45) into the equality constraint of (44), we have
$y_{1}$ $=Y_{\textnormal{F}}H_{1}^{\dagger}\begin{bmatrix}u_{\textnormal{ini}}\\ y_{\textnormal{ini}}+\sigma_{y_{1}}\\ u_{1}\end{bmatrix}$ (47)
$=(M+M^{\bot})H_{1}^{\dagger}\begin{bmatrix}u_{\textnormal{ini}}\\ y_{\textnormal{ini}}+\sigma_{y_{1}}\\ u_{1}\end{bmatrix}$
$=MH_{1}^{\dagger}\begin{bmatrix}u_{\textnormal{ini}}\\ y_{\textnormal{ini}}+\sigma_{y_{1}}\\ u_{1}\end{bmatrix}$.
Upon defining $g_{1}=H_{1}^{\dagger}\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y_{1}},u_{1})$, we have $y_{1}=Mg_{1}$ from (47). We then substitute $g_{1}$ into the equality constraint of (21), leading to
$\begin{bmatrix}H_{1}\\ M\end{bmatrix}g_{1}=\begin{bmatrix}H_{1}H_{1}^{\dagger}\begin{bmatrix}u_{\textnormal{ini}}\\ y_{\textnormal{ini}}+\sigma_{y_{1}}\\ u_{1}\end{bmatrix}\\ Mg_{1}\end{bmatrix}=\begin{bmatrix}u_{\textnormal{ini}}\\ y_{\textnormal{ini}}+\sigma_{y_{1}}\\ u_{1}\\ y_{1}\end{bmatrix}$,
where we have used the fact $H_{1}H_{1}^{\dagger}=I$ since $H_{1}$ has full row rank. This means that $u_{1},y_{1},\sigma_{y_{1}},g_{1}$ is a feasible solution for (21). This completes our proof.
B.2 Proof of Proposition 1
As (19) and (21) are equivalent directly from construction, we here mainly prove the equivalence of (21) and (22). It is obvious that (21) and (22) have the same objective function and it only contains decision variables $u,y$ and $\sigma_{y}$. Thus, we show that (21) and (22) provide the same unique optimal solution $u^{*},y^{*}$ and $\sigma_{y}^{*}$ by proving:
  1. 1.
    Feasible regions of $u,y,\sigma_{y}$ are the same for (21) and (22): if $u,y,\sigma_{y},g$ is feasible to (22), then the same $u,y,\sigma_{y},g$ is also feasible for (21). Conversely, given any feasible solution $u,y,\sigma_{y},g$ to (21), there exists a vector $\tilde{g}$ such that $u,y,\sigma_{y},\tilde{g}$ is feasible to (22).
  2. 2.
    The optimal solution of $u,y,\sigma_{y}$ is unique for (21).
We assume $u_{1},y_{1},\sigma_{y_{1}},g_{1}$ is a feasible solution for (22). Substituting the orthogonal decomposition (45) of $Y_{\textnormal{F}}$ into (22a), we have
$H_{\textnormal{d}}g_{1}=\!\begin{bmatrix}U_{\textnormal{P}}\\ Y_{\textnormal{P}}\\ U_{\textnormal{F}}\\ M+Y_{\textnormal{F}}(I-\Pi_{1})\end{bmatrix}g_{1}\!=H_{\textnormal{s}}^{*}g_{1}\!=\!\begin{bmatrix}u_{\textnormal{ini}}\\ y_{\textnormal{ini}}+\sigma_{y_{1}}\\ u_{1}\\ y_{1}\end{bmatrix}$,
[…]
We next assume $u_{1},y_{1},\sigma_{y_{1}}$ and $g_{1}$ is a feasible solution for (21). We define $\tilde{g}_{1}=H_{1}^{\dagger}\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y_{1}},u_{1})$, which satisfies $y_{1}=Y_{\textnormal{F}}\tilde{g}_{1}$ from (46). We first verify that $\tilde{g}_{1}$, together with $u_{1},y_{1},\sigma_{y_{1}}$, satisfies (22a):
$H_{\textnormal{d}}\tilde{g}_{1}=\begin{bmatrix}H_{1}\\ Y_{\textnormal{F}}\end{bmatrix}H_{1}^{\dagger}\begin{bmatrix}u_{\textnormal{ini}}\\ y_{\textnormal{ini}}+\sigma_{y_{1}}\\ u_{1}\end{bmatrix}=\begin{bmatrix}u_{\textnormal{ini}}\\ y_{\textnormal{ini}}+\sigma_{y_{1}}\\ u_{1}\\ y_{1}\end{bmatrix}$.
For the satisfaction of (22b), since $\tilde{g}_{1}$ is in the range space of $H_{1}^{\dagger}$ and $\Pi_{1}$ is the orthogonal projector onto the row space of $H_{1}$, we have $\Pi_{1}\tilde{g}_{1}=\tilde{g}_{1}$ (the range space of $H_{1}^{\dagger}$ and row space of $H_{1}$ are equivalent), which implies $\|(I-\Pi_{1})\tilde{g}_{1}\|_{2}=\|\tilde{g}_{1}-\tilde{g}_{1}\|_{2}=0$. Thus, $u_{1},y_{1},\sigma_{y_{1}}$ and $\tilde{g}_{1}$ is a feasible solution for (22).
The uniqueness of the optimal solution $u^{*},y^{*}$ and $\sigma_{y}^{*}$ for (21) basically comes from strong convexity. For notational simplicity, we define $x_{\textnormal{d}}=\textnormal{col}(u,y,\sigma_{y},g)$ and $f_{\textnormal{d}}(x_{\textnormal{d}})=\|y\|_{Q}^{2}+\|u\|_{R}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2}$ as the decision variable and objective function of (21), respectively. Suppose that $x_{\textnormal{d}_{1}}$ and $x_{\textnormal{d}_{2}}$ are two optimal solutions with different $u,y$ or $\sigma_{y}$. Let the optimal value be $f_{\textnormal{d}}^{*}$. We then construct a convex combination $x_{\textnormal{d}_{3}}=\alpha x_{\textnormal{d}_{1}}+(1-\alpha)x_{\textnormal{d}_{2}}$ where $0<\alpha<1$. This new point $x_{\textnormal{d}_{3}}$ is also feasible as
$\begin{bmatrix}U_{\textnormal{P}}\\ Y_{\textnormal{P}}\\ U_{\textnormal{F}}\\ M\end{bmatrix}g_{3}$ $=\begin{bmatrix}U_{\textnormal{P}}\\ Y_{\textnormal{P}}\\ U_{\textnormal{F}}\\ M\end{bmatrix}(\alpha g_{1}+(1-\alpha)g_{2})$
$=\begin{bmatrix}u_{\textnormal{ini}}\\ y_{\textnormal{ini}}+\alpha\sigma_{y_{1}}+(1-\alpha)\sigma_{y_{2}}\\ \alpha u_{1}+(1-\alpha)u_{2}\\ \alpha y_{1}+(1-\alpha)y_{2}\end{bmatrix}=\begin{bmatrix}u_{\textnormal{ini}}\\ y_{\textnormal{ini}}+\sigma_{y_{3}}\\ u_{3}\\ y_{3}\end{bmatrix}$.
It is obvious that $f_{\textnormal{d}}(\cdot)$ is a strongly convex function with respect to $u,y$ and $\sigma_{y}$ and its value is not affected by $g$. Thus, we have $f_{\textnormal{d}}(x_{\textnormal{d}_{3}})=f_{\textnormal{d}}(\alpha x_{\textnormal{d}_{1}}+(1-\alpha)x_{\textnormal{d}_{2}})<\alpha f_{\textnormal{d}}(x_{\textnormal{d}_{1}})+(1-\alpha)f_{\textnormal{d}}(x_{\textnormal{d}_{2}})=f_{\textnormal{d}}^{*},$ which contradicts our assumption. The optimal solution to (21) is thus unique. This completes our proof.
B.3 Proof of Proposition 2
Similar to the proof in the Appendix B.2, we here mainly prove feasible regions of $u,y,\sigma_{y}$ are equivalent for (26) and (28). Then, with the same objective function, their optimal solutions are also the same. The proof for the uniqueness of the optimal solution utilizes the strong convexity of the objective function of (26) following the same procedure in Appendix B.2. We omit the proof for the uniqueness here. We recall that we have $H_{1}=\textnormal{col}(U_{\textnormal{P}},Y_{\textnormal{P}},U_{\textnormal{F}})=\begin{bmatrix}L_{11}&\mathbb{0}\\ L_{21}&L_{22}\end{bmatrix}\begin{bmatrix}Q_{1}\\ Q_{2}\end{bmatrix}$ from the LQ decomposition of $H_{\textnormal{d}}$. As we assume $H_{\textnormal{d}}$ has full row rank, the matrix $\begin{bmatrix}L_{11}&\mathbb{0}\\ L_{21}&L_{22}\end{bmatrix}$ is invertible.
Let $u_{1},y_{1},\sigma_{y_{1}}$ and $g_{1}$ be a set of feasible solution for (28). Substituting (28b) into (28a) leads to
$\hat{H}g_{1}\!=\!\begin{bmatrix}L_{11}&\mathbb{0}&\mathbb{0}&\mathbb{0}\\ L_{21}&L_{22}&\mathbb{0}&\mathbb{0}\\ L_{31}&L_{32}^{*}&L_{33}&L_{32}^{\prime}\end{bmatrix}\!\begin{bmatrix}Q_{1}g_{1}\\ Q_{2}g_{1}\\ \mathbb{0}\\ \mathbb{0}\end{bmatrix}\!=\!H_{\textnormal{sc}}g_{1}\!=\!\begin{bmatrix}u_{\textnormal{ini}}\\ y_{\textnormal{ini}}+\sigma_{y_{1}}\\ u_{1}\\ y_{1}\end{bmatrix}$,
thus, $u_{1},y_{1},\sigma_{y_{1}}$ and $g_{1}$ is also a feasible solution for (26).
We next choose a feasible solution $u_{1},y_{1},\sigma_{y_{1}}$ and $g_{1}$ of (26). We can construct $\tilde{g}_{1}$ as
$\tilde{g}_{1}$ $=H_{1}^{\dagger}\begin{bmatrix}u_{\textnormal{ini}}\\ y_{\textnormal{ini}}+\sigma_{y_{1}}\\ u_{1}\end{bmatrix}$ (48)
$=\begin{bmatrix}Q_{1}^{{\mathsf{T}}}&Q_{2}^{{\mathsf{T}}}\end{bmatrix}\begin{bmatrix}L_{11}&\mathbb{0}\\ L_{21}&L_{22}\end{bmatrix}^{-1}\begin{bmatrix}u_{\textnormal{ini}}\\ y_{\textnormal{ini}}+\sigma_{y_{1}}\\ u_{1}\end{bmatrix}$
and $g_{1}$ can be represented as
$g_{1}=\tilde{g}_{1}+\hat{g}$
where $\hat{g}\in\textrm{null}(H_{1})=\textrm{null}(\textnormal{col}(Q_{1},Q_{2}))$ and the equality is derived from the matrix $\begin{bmatrix}L_{11}&\mathbb{0}\\ L_{21}&L_{22}\end{bmatrix}$ is invertible. Thus, the $\tilde{g}_{1}$ satisfies
[…]
$=\begin{bmatrix}L_{31}&L_{32}^{*}\end{bmatrix}\begin{bmatrix}Q_{1}\\ Q_{2}\end{bmatrix}g_{1}=y_{1}.$
Since $\textnormal{col}(Q_{1},Q_{2},Q_{3},Q^{*})$ has orthonormal rows, we have
$Q_{\textnormal{c}}\tilde{g}_{1}=\underbrace{\begin{bmatrix}Q_{3}\\ Q^{*}\end{bmatrix}\begin{bmatrix}Q_{1}^{{\mathsf{T}}}&Q_{2}^{{\mathsf{T}}}\end{bmatrix}}_{=\mathbb{0}}\begin{bmatrix}L_{11}&\mathbb{0}\\ L_{21}&L_{22}\end{bmatrix}^{-1}\begin{bmatrix}u_{\textnormal{ini}}\\ y_{\textnormal{ini}}+\sigma_{y_{1}}\\ u_{1}\end{bmatrix}=\mathbb{0}$
and (28b) is naturally satisfied. We can further substitute $\tilde{g}_{1}$ to (28a), leading to
$\begin{bmatrix}L_{11}&\mathbb{0}&\mathbb{0}&\mathbb{0}\\ L_{21}&L_{22}&\mathbb{0}&\mathbb{0}\\ L_{31}&L_{32}^{*}&L_{33}&L_{32}^{\prime}\end{bmatrix}\begin{bmatrix}Q_{1}\\ Q_{2}\\ Q_{3}\\ Q^{*}\end{bmatrix}\tilde{g}_{1}$ (49)
$=$ $\begin{bmatrix}u_{\textnormal{ini}}\\ y_{\textnormal{ini}}+\sigma_{y_{1}}\\ u_{1}\\ \begin{bmatrix}L_{31}&L_{32}^{*}\end{bmatrix}\begin{bmatrix}Q_{1}\\ Q_{2}\end{bmatrix}\tilde{g}_{1}\\ \end{bmatrix}=\begin{bmatrix}u_{\textnormal{ini}}\\ y_{\textnormal{ini}}+\sigma_{y_{1}}\\ u_{1}\\ y_{1}\\ \end{bmatrix}$.
Thus, $u_{1},y_{1},\sigma_{y_{1}}$ and $\tilde{g}_{1}$ is a feasible solution for (28).
B.4 Proof of Corollaries 1 and 2
We prove Corollaries 1 and 2 via showing both (22) and (28) can be represented in the form of (17). Thus, Corollaries 1 and 2 are actually special cases of Theorem 1.
For changing (22) to the same form as (17), we can let
$x_{1}:=\textnormal{col}(\sigma_{y},u,y),\quad x_{2}:=g,\quad\mathcal{X}:=\Gamma\times\mathcal{U}\times\mathcal{Y},$
$A_{1}:=\textnormal{col}(\mathbb{0},I),\quad A_{2}:=-H_{\textnormal{d}},\quad b:=-\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}},\mathbb{0}),$
$M:=\mathrm{diag}(\lambda_{y}I,R,Q),\quad D:=(I-\Pi_{1}).$
Similarly, we transform (28) to the form of (17) by letting
$x_{1}:=\textnormal{col}(\sigma_{y},u,y),\quad x_{2}:=g,\quad\mathcal{X}:=\Gamma\times\mathcal{U}\times\mathcal{Y},$
$A_{1}:=\textnormal{col}(\mathbb{0},I),\quad A_{2}:=-\hat{H},\quad b:=-\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}},\mathbb{0}),$
$M:=\mathrm{diag}(\lambda_{y}I,R,Q),\quad D:=Q_{\textnormal{c}}.$
We note that the only differences between transformations of (22) and (28) are the choices of $D$ and $A_{2}$.
As $\Gamma,\mathcal{U},\mathcal{Y}$ are convex and compact sets, $\mathcal{X}$ defined above for (22) and (28) are convex and compact. Furthermore, a quadratic function is locally Lipschitz continuous over a compact set. Thus, all conditions in Theorem 1 are satisfied and that completes the proof.
B.5 Proof of Theorem 4
The key idea for the proof of Theorem 4 is that KKT conditions of (32) and (33) have similar forms. We first illustrate that the primal and dual optimal solutions of (32) satisfy its KKT condition. Then, we present that they also satisfy the KKT condition for (33) under a specific choice of $\lambda_{g}$. That is sufficient to ensure the optimal solution of (32) is also that of (33).
We denote $x:=\textnormal{col}(\sigma_{y},u,y)$, $v:=\textnormal{col}(u_{\textnormal{ini}},y_{\textnormal{ini}}+\sigma_{y},u,y)$ and we define $f(x):=\|y\|_{Q}^{2}+\|u\|_{R}^{2}+\lambda_{y}\|\sigma_{y}\|_{2}^{2}$. Let $x^{*},g^{*}$ be the primal optimal solution of (32) and $\mu_{1}^{*},\mu_{2}^{*}$ be the dual optimal solutions for the inequality constraint and equality constraint of (32) respectively. Since (32) satisfies Slater’s condition, $x^{*},g^{*},\mu_{1}^{*}$ and $\mu_{2}^{*}$ satisfy the KKT condition and have the follow properties
$0\in\partial(f(x)+\mu_{1}^{*}(\|g\|_{1}-\alpha_{\textnormal{c}})+\mu_{2}^{*{{\mathsf{T}}}}(Hg-v^{*}))$ (50)
$\text{at}\ (x,g)=(x^{*},g^{*})\ \text{and}\ Hg^{*}=v^{*}.$
We can choose $\lambda_{g}^{*}=\mu_{1}^{*}$ and then the KKT conditions for the primal solution $\bar{x},\bar{g}$ and dual solution $\bar{\mu}$ of (33) are
$0\in\partial(f(x)+\mu_{1}^{*}\|g\|_{1}+\bar{\mu}^{{\mathsf{T}}}(Hg-v)),$ (51)
$\text{at}\ (x,g)=(\bar{x},\bar{g})\ \text{and}\ H\bar{g}=\bar{v}.$
We can let $\bar{x}=x^{*},\bar{g}=g^{*}$ and $\bar{\mu}=\mu_{2}^{*}$. Then, the KKT condition (51) holds as it becomes equivalent to (50). Thus, $(\sigma_{y}^{*},u^{*},y^{*},g^{*})$ (i.e., $(x^{*},g^{*})$) is also an optimal solution for (33) as the KKT condition is sufficient to guarantee optimality.
B.6 Proof of Proposition 5
The key idea for the proof of the Proposition 5 is that we can first partition both $H_{y}$ and $\tilde{H}_{y}$ into the part in the row space of $H_{u}$ and the part in the null space of $H_{u}$, i.e., $H_{y}=H_{1}+H_{2}$, $\tilde{H}_{y}=\tilde{H}_{1}+\tilde{H}_{2}$, where $\mathrm{row}(H_{1}),\mathrm{row}(\tilde{H}_{1})\subseteq\mathrm{row}(H_{u})$ and $\mathrm{row}(H_{2}),\mathrm{row}(\tilde{H}_{2})\subseteq\mathrm{null}(H_{u})$. Then, we illustrate that the optimal solution for $\tilde{H}_{1}$ is $H_{1}$ and thus the problem becomes approximately $H_{2}$ with a low rank matrix $\tilde{H}_{2}$ which admits an analytical solution via SVD.
We can equivalently formulate the optimization problem (35) as
$\min_{\tilde{H}_{1},\tilde{H}_{2}}$ $\|H_{1}+H_{2}-\tilde{H}_{1}-\tilde{H}_{2}\|_{F}$ (52)
$\mathrm{subject~to}$ $\textrm{rank}(\tilde{H}_{2})=n,$
$\mathrm{row}(\tilde{H}_{1})\subseteq\mathrm{row}(H_{u}),$
$\mathrm{row}(\tilde{H}_{2})\subseteq\mathrm{null}(H_{u}).$
The equivalence of (35) and (52) comes from
  1. 1.
    For any feasible solution $\tilde{H}_{y}$ of (35), we can let $\tilde{H}_{1}=\tilde{H}_{y}\Pi_{2}$ and $\tilde{H}_{w}=\tilde{H}_{y}(I-\Pi_{2})$, which satisfies constraints for (52) and provides the same value for the objective function.
  2. 2.
    Conversely, for any feasible solution $\tilde{H}_{1}$ nad $\tilde{H}_{2}$ of (52), we can let $\tilde{H}_{y}=\tilde{H}_{1}+\tilde{H}_{2}$ which is feasible for (35) with the same objective function value.
We can expand the objective function in (52) as
$\|H_{1}+H_{2}-\tilde{H}_{1}-\tilde{H}_{2}\|_{F}$
$=$ $\ \mathrm{tr}((H_{1}+H_{2}-\tilde{H}_{1}-\tilde{H}_{2})(H_{1}+H_{2}-\tilde{H}_{1}-\tilde{H}_{2})^{{\mathsf{T}}})$
$=$ $\ \mathrm{tr}((H_{1}-\tilde{H}_{1})(H_{1}-\tilde{H}_{1})^{{\mathsf{T}}})+\mathrm{tr}((H_{2}-\tilde{H}_{2})(H_{2}-\tilde{H}_{2})^{{\mathsf{T}}})$
$\ +2\mathrm{tr}((H_{1}-\tilde{H}_{1})(H_{2}-\tilde{H}_{2})^{{\mathsf{T}}}))$
$=$ $\ \|H_{1}-\tilde{H}_{1}\|_{F}+\|H_{2}-\tilde{H}_{2}\|_{F},$
where the third equality is derived from $\mathrm{row}(H_{1}),\mathrm{row}(\tilde{H}_{1})\subseteq\mathrm{row}(H_{u})$ and $\mathrm{row}(H_{2}),\mathrm{row}(\tilde{H}_{2})\subseteq\mathrm{null}(H_{u})$. Thus, it is obvious that the optimal solution of $\tilde{H}_{1}$ is $H_{1}$ and we can simplify (52) as
$\min_{\tilde{H}_{2}}$ $\|H_{2}-\tilde{H}_{2}\|_{F}$ (53)
$\mathrm{subject~to}$ $\textrm{rank}(\tilde{H}_{2})=n,$
$\mathrm{row}(\tilde{H}_{2})\subseteq\mathrm{null}(H_{u}),$
which admits an analytical solution from the SVD that is $\tilde{H}_{2}^{*}=\sum_{i=1}^{n}\bar{\sigma}_{i}\bar{u}_{i}\bar{v}_{i}^{{\mathsf{T}}}$ where $H_{2}=\sum_{i=1}^{pL}\bar{\sigma}_{i}\bar{u}_{i}\bar{v}_{i}^{{\mathsf{T}}}$. We note that $\mathrm{row}(\tilde{H}_{2}^{*})\subseteq\mathrm{row}(H_{2})\subseteq\mathrm{null}(H_{u})$. That completes the proof.